A Complete Bibliography of Publications in Reviews of Modern Physics (1990–1999)

Nelson H. F. Beebe University of Utah Department of Mathematics, 110 LCB 155 S 1400 E RM 233 Salt Lake City, UT 84112-0090 USA

Tel: +1 801 581 5254 FAX: +1 801 581 4148

       E-mail: beebe@math.utah.edu, beebe@acm.org,
                  beebe@computer.org (Internet)
     WWW URL: https://www.math.utah.edu/~beebe/

02 November 2023 Version 1.05

Title word cross-reference

(151 ≤ A ≤ 193) [JSSJ90]. (A ≥ 221) [JSSJ90]. (e, 2e) [CMD94b, CMD94a]. (Z < 6) [Ree94]. 1/2 [Man91]. 144 < A < 194 [JSH^{+}98]. 17 [Fra95]. ^{2} [BK96a]. ^{3} [Lee97, Lee98, Osh97, SO95]. _{14} [Her91a]. _{2} [Her91a]. _{3}^{+} [Oka92]. 60 [Kro97, PS96]. A [JSSJ90]. d [SR95]. d_{x^{2}−y^{2}} [Van95]. g [HK99]. μ [Ama97]. n [LR97]. N_{c} [Coh96]. x [BCKK92]. Z [MMRT99].

-body [LR97]. -keV [Fra95]. -wave [SR95].

1996 [BoP96].

61 [JP90]. 69 [Lee98].

70 [Gre00]. 71 [KKL^{+}99a, Wei99a]. absorption [Swa95]. Accelerators [PB99, HK91]. accretion [BH98]. Acknowledgments [FKT91]. acoustic [YB95]. action [HN95, LSST99]. action-at-a-distance [HN95]. Addendum [JP90, KM91a]. adiabatic [BT99]. adsorption [Eva93]. Adventures [CG99]. after [For95, HTB90]. Alaska [SO95]. algebra [FS93]. algebras [KM91b]. algorithm [EJ96]. alkali [BPS90, BPS91]. alloys [FAG^{+}94]. among [BTS98]. analysis [ABST93, Ari99, BBF^{+}94, Gil98]. Analytical [GL91]. andapplication [Hir97]. Anderson [BK94, Pac98]. andinternal [LR97]. anharmonic [BHK98]. Anisotropic [CT94, CG99]. annealing [Sah93]. annihilation [Ams98]. Anomalous [HK99, CT94]. antiferromagnet [Man91]. antiferromagnetism [FAG^{+}94]. antiproton [Ams98]. antiprotons [EH99]. appearance [Fra95]. Application [Gre98, Man91, Gre00]. applications [HHI^{+}93, Hen99, KM91b, Rob91, YB95, ZFG90]. Approach [Moo93, EPV95, PK98, Ram91, Res94, Sha94]. approaches [Bra93a, BT99]. approximants [GK93b, GK93a]. arbitrary [Cah96]. artificial [Win92]. aspects [Her91a, RT92, ST92, dH93]. Astrometry [SFJ98]. astronomy [HSFP99, Ost93, Rob92]. astrophysical [AL97]. Astrophysics [Fie99]. asymmetric [Rau92]. asymmetry [BN96]. atlas [Bro95]. Atom [WPW99]. Atomic [BC94, DDG^{+}99, BN96, Kle99b, LBNZ98, Sap98]. atoms [AGLK96, BTS98, BR99, Bra93b, CT98, Phi98, RKMG92, Spr91, SP98]. attention [Jul91]. attractive [MRR90]. automata [Sah93]. avalanches [Cha93]. axial [RMRW91].

B [Her91a]. background [PW99]. baked [SO95]. band [MRR90, Nen91, Van92]. bang [ST98]. Barrel [Ams98]. Barrier [LM94]. baryogenesis [Tro99]. baryon [Coh96]. based [RWF97]. baseline [Rob91]. bases [Ari99]. basic [KM94]. basis [Fis98]. beam [Sis93]. beams [CSB98, Win92]. behavior [CCDR^{+}93, Kol99]. Bell [Mer16, Mer93]. beyond [BBL96, LLM92]. bifurcation [Cra91]. Big [ST98]. Big-bang [ST98]. Binary [Tay94, Hul94]. Biological [FWA99, KM94]. birth [Rob92]. Black [HT99b]. body [FS93, Gut98, LR97, Sap98, Wir93]. Boltzmann [Bav91]. Bose [DGPS99]. Boson [KM91b, Car90]. boson-exchange [Car90]. bosonic [CD94]. bottom [RB95]. Boundary [Fre90, Fre91]. bounds [PF95]. Brain [Hop99, HHI^{+}93]. Brane [GK99]. Breaking [For95]. breakup [Egg97]. Bremsstrahlung [Kle99a]. brown [BL93]. Built [Kad99]. bulk [Spr91].

C [Kro97, PS96]. calculations [PTA^{+}92]. cavity [Vav94]. cells [RB98]. cellular [Sah93]. central [Leb99]. century [HS99, Koh99a, Sha99, Kle99b]. chaos [BBF^{+}94, Moo93]. Chaotic [Lis99, NF95, ABST93, Gil98]. characterization [YB95]. charged [DROM92, Pau90]. charm [RB95]. Chemical [Her99, Pop99]. chemistry [Mar93]. Chiral [Coh96, HKL99]. chirality [HKL99]. chromium [FAG^{+}94]. chromodynamics [Coh96, Kle92, MO96]. Classical [LBNZ98, BO96, dVvCL98, vRN99]. clouds [Her99]. cluster [Jen99]. clusters [Bra93a, Her99, dH93]. Coherent [BHK98, BTS98, ZFG90]. cold [WBZJ99]. collapse [Rob97]. collective [Fan92]. collider [AC99, CCF^{+}99]. collisionless [OC99]. collisions [RKMG92, WBZJ99]. Colloquia [Pin92]. color [Lif97]. common [AM98, Fan92]. communication [CD94]. communications [BL99, HW96]. Comparisons [Fri91]. complex [AGLK96, KM94]. compound [MBW99]. compounds [Deg99]. computation [EJ96, Hop99]. concept [Hir97]. condensation [Cve99, DGPS99]. condensations [Mee96]. condensed [Cep95, Gol96, Koh99a, Wit99]. conditions [Fre90, Fre91]. Conductance [IL99]. conduction [TBW99]. conduction-electron [TBW99]. conduction-electron/interface [TBW99]. confined [Gen95]. confinement [Haw98]. conical [Yar96]. conjecture [Cur97]. conjugate [PTA^{+}92]. consequence [SR95]. Consistent [Omn92, Bra93a]. Continuous [SGCS97, Das99]. continuum [Sah93]. Controlled [Win92]. controversy [LSST99]. conversion [Swa95]. Cooling [Pet92, Phi98, WPW99]. cooperative [Eva93]. copper [KBSE98]. corrals [Her99]. Correlated [Dag94, GKKR96]. Correlation [Shu93, KS96, SM94]. correlations [BPP93, PSdH97]. Cosmic [Cro99, PW99]. Cosmology [HN95, TT99]. coupled [BO96]. Coupling [HK91, KGJ^{+}99, LBM94]. CP [WW93]. critical [Sta99]. cross [A^{+}98, LLL93, SS96]. crystal [SB99, Ams98]. crystalline [GK93a, GK93b, MBB^{+}92, Res94]. crystals [DO99, Lif97, LBM94, Mer92c, RMRW91, Mer92a, Mer92b, Mer94]. cubic [Mer92a, Mer92b, Mer92c, Mer94]. cuprous [Man91]. current [CCNR91, FS93]. Cyclotron [Swa95].

dark [Sad99]. data [ABST93]. Dawn [Cur97]. Decay [SS96]. decays [BBL96, RB95, RW93]. Deciphering [Sad99]. Deep [Fri91, FKT91, Ken91, Tay91, BCKK92]. deep-inelastic [BCKK92]. defects [MBB^{+}92]. deformed [JSSJ90]. dense [Her91b, Ich93]. density [FAG^{+}94, Grü94, Koh99b]. dependent [BPS90, BPS91, DDLÖ94]. deposition [Jen99]. description [Mor98]. detection [RB98]. detectors [PB99]. deuterium [Haw98]. deuterium-tritium [Haw98]. deuterons [Wir93]. development [Shu95]. developments [OT90]. devices [Gre98, Gre00, KKL^{+}99a, KKL^{+}99b]. Diabolical [Yar96]. Diagnostics [Gen95]. diagram [KS96, TSU97]. dielectrics [Res94, ST92]. diffraction [CT94]. diffusion [BP92, BPW95, MG98, SP98, vRN99]. diffusion-limited [MG98]. Digest [BoP96]. dimension [KS96]. dimensional [Tsu99, TSU97]. dimensions [GKKR96]. Dipole [VG90, BS91, BS92]. Diquarks [APE^{+}93]. direct [WW93]. disappearance [Fra95]. Discovering [Sma97]. Discovery [Osh97, CF97, Hul94]. Discrete [Bra93b]. disks [BH98]. disorder [Tsu99]. disordered [LBM94]. dissipative [LCD^{+}87, LCD^{+}95, PK98]. distance [HN95]. distributions [RKMG92]. Do [GL91]. doped [Van92]. Drift [Hor99]. drive [CCNR91]. driven [Fre90, Fre91]. duality [Pol96]. due [Kle99a]. DUMAND [Rob92]. dwarfs [BL93]. dying [AL97]. Dynamic [CA99, Iac93, Bis90]. Dynamical [GKKR96, NYO94, Gil98]. Dynamics [CLS94, KM89, LCD^{+}87, Nen91, CS98, DDLÖ94, Egg97, GK99, Grü94, Mee96, PTA^{+}92, PK98, KM91a, LCD^{+}95].

Early [Jul91, Pei93, Shu95, Tay91]. Earth [Gut98, Win92]. earthquake [CLS94]. edge [OT90]. Editorial [Pin92]. effect [Huc95, Ric97, STG99, Sto99]. effective [Nen91]. effects [Man99, SO95]. eighteen [Cah96]. Einstein [DGPS99]. Elastic [Vav94, LLL93, LBM94]. electric [BS91, BS92, Bis90, Nen91, VG90]. electrical [BPS90, BPS91]. electrodynamic [Deg99]. electrodynamics [HN95]. Electromagnetic [Pau90]. Electron [BPP93, Her91b, Mar93, RKMG92, BS91, CG99, CCNR91, DCCL91, Deg99, HK99, Kas92, PR92, SM94, TBW99, Tsu99, VM95, Win92, BS92]. Electron-scattering [BPP93]. electron/interface [TBW99]. Electronic [Cha93, Koh99b, Ari99, Goe99, SM90]. electrons [Dag94, DDLÖ94, Nen91, Ram91]. electrophotography [PS93]. electroproduction [BK96a]. Electroweak [Ros92, Tro99, LLM92, MMRT99, Ros94]. element [BPW95]. Elementary [Fit99]. elements [Ree94, WIP^{+}97]. Emden [Bav91]. emission [Swa95]. end [HS99]. energetic [Cro99]. energies [HSFP99]. energy [CSB98, PTA^{+}92, PF95, Rob92, RKMG92, Vav94]. enhanced [BH98]. enters [ST98]. environmental [Kle99a]. equation [Bav91, Pac98]. equations [MMTB96]. Equilibrium [Bav91, DCCL91, CH93, DO99, Fre90, Fre91]. era [ST98]. Erratum [BPS91, BS92, CMD94b, Fre91, GK93a, Gre00, KKL^{+}99a, Lee98, LCD^{+}95, Mer92a, Mer92b, Mer94, Mer16, Ros94, Rup96, She95, Wei99a, Wu93]. essay [Koh99a]. everyday [Cah96]. Evidence [Van95]. evolution [AL97, ODB97]. exchange [Car90, WH97]. excitations [Ost92, Qui97]. exist [GL91]. expansion [CLvdB^{+}98]. Experiment [Zei99, Cur97, Mar93]. experimental [BBF^{+}94, BO96, DROM92, dH93]. Experiments [Deh90, Jen99, Ken91, Ram90, WBZJ99, Haw98, Kir99, LLM92, SFJ98, Win92]. exponents [KS96]. external [Bra93b]. extraordinary [Lee97, Lee98].

factoring [EJ96]. factors [Kle99a]. fall [DROM92]. far [Fre90, Fre91]. fate [AL97]. faults [CLS94]. Fe [Her91a]. femtosecond [Blo99]. Fermi [Spr91]. fermion [Ama97, GKKR96, PSdH97]. fermions [KS96, Sha94]. ferroelectricity [VG90]. ferroelectrics [BHK98]. few [CS98]. few-nucleon [CS98]. field [For92, GKKR96, MG98, SM94, Vav94, Wil99]. fields [Bis90, Bra93b, LR97, Nen91, Ram90, RT92, ST92, Tsu99]. fifty [HTB90]. films [CCDR^{+}93, ODB97]. fission [Til99]. fixed [EPV95]. fixed-scale [EPV95]. Flow [Sah93, DCCL91, RZ94]. flows [Egg97, Kad99]. fluctuation [KG99, Rup95, Rup96]. fluctuation-induced [KG99]. Fluid [Sre99, Mor98]. focal [Jul91]. focusing [TBW99]. forbidden [CT94]. force [BO96, SGB^{+}95]. forces [KG99]. formation [CH93, GL99, KM94]. formulation [Fis98]. Forty [EH99, WIP^{+}97]. foundations [Zei99]. fractal [EPV95, NYO94]. fractals [Sah93]. Fractional [Lau99, For92, STG99, Sto99]. free [CCNR91, Egg97]. free-electron-laser [CCNR91]. free-surface [Egg97]. friction [KG99]. fullerenes [Cur97, Sma97]. fulleride [PS96]. fullerides [Gun97]. functional [Bav91]. functionals [Koh99b]. functions [Koh99b, Shu93]. fusion [A^{+}98, BT98, DD95, Fow99, Ich93, She94, She95].

galaxies [HT99a]. Gamma [HSFP99, BS97a]. Gamma-ray [HSFP99, BS97a]. gas [RZ94, Tsu99, WH97]. gases [Cha93, DGPS99, Her91b, JNB96]. Gauge [FS93, LR97, GK99]. general [SS94]. Generalized [PF95, Bav91]. geodesy [SFJ98]. geometric [Mea92, Res94, Wit98]. geometry [Rup95, Rup96]. Geophysical [Rob91]. glass [VG90, Wei93]. glimpse [Wei93]. global [LLM92]. gradients [PTA^{+}92]. grand [Bro95]. Granular [JNB96, dG99, Kad99, SR95]. Gravitational [Sut99, Wei99b, Wei99a]. gravity [DROM92, Tay94]. gravomagnetic [LBNZ98]. ground [For95, RWF97, TSU97]. ground-based [RWF97]. ground-state [TSU97]. group [Fis98, Sha94]. groups [Mer92a, Mer92b, Mer92c, Mer94, RMRW91]. Growth [Jen99, EPV95, MMTB96].

H [Oka92]. hadron [BCKK92]. hadrons [RB95]. Hall [Huc95, STG99, Sto99]. halo [Rii94]. Hamiltonian [Mor98]. Hamiltonians [Nen91]. Handbook [SSC^{+}95]. hardware [For95]. harmonic [BO96]. Heat [JP89, JP90]. heating [CCNR91]. heavily [Van92]. Heavy [Ama97, BPW95, Deg99]. heavy-electron [Deg99]. heavy-element [BPW95]. Heavy-fermion [Ama97]. Heisenberg [Man91]. helium [BP92, BPW95, Cep95, NF95, Pac98]. HERA [AC99]. Hidden [Mer93, Mer16]. High [KKL^{+}99b, LLM92, BFG^{+}94, CSB98, Dag94, Gen95, HSFP99, OC99, RT92, Rob92, SR95, Van95, YMS96, KKL^{+}99a]. high-energy [CSB98, Rob92]. High-precision [LLM92]. high-temperature [BFG^{+}94, Dag94, Gen95, OC99, SR95, Van95, YMS96]. High-transition-temperature [KKL^{+}99b, KKL^{+}99a]. historical [OT90]. History [Das99, Kle99b, Rob92]. holes [HT99b]. human [HHI^{+}93]. hydrodynamic [NF95]. Hydrogen [MBB^{+}92, kMH94, Ram90]. hypotheses [SS94]. hysteresis [CA99].

ices [Joh96]. icosahedral [Mer92a, Mer92b, Mer92c, Mer94]. ideal [Mor98]. ideas [Kad99]. II [For95]. imaging [Cha93, Hen99]. Immunology [PW97]. impedance [HK91]. implications [Kir99]. improvements [Spr91]. Improving [RWF97]. Independent [PSdH97]. indeterminacy [BU95]. induced [KG99]. industry [BL99]. inelastic [BCKK92, Fri91, FKT91, Ken91, Tay91]. infinite [GKKR96, Seg91]. infinitesimal [Seg91]. inflaton [LLK^{+}97]. infrared [Oka92]. initio [PTA^{+}92]. Insights [Wit99]. inspired [Kad99]. Instabilities [Nag92, Moo93]. Instability [BH98]. Instantons [SS98]. instrumentation [HHI^{+}93]. insulator [IFT98]. integer [Huc95]. integrable [KM89, KM91a]. integrals [Cep95]. intense [DCCL91, Tsu99]. Intensity [BGJ90]. interacting [KS96, Sha94]. interaction [BT99, LM94, MMRT99, Tsu99]. interactions [MRR90, MBB^{+}92, TBW99]. interface [TBW99]. interfaces [RZ94]. interference [Gre98, Gre00, KKL^{+}99a, KKL^{+}99b, Man99]. interferometry [BGJ90, Rob91, SFJ98]. Interplay [Tsu99]. interpretations [Omn92]. intersections [Yar96]. interstellar [BKS95, HT99a]. Intrinsic [BN96, Her91a, JSSJ90]. Introduction [Cra91, Wit98]. invariance [FS93, MMTB96]. invention [RHH99]. ion [BC94]. ionization [Sis93]. ionizing [Cha93]. isolated [Deh90]. isospin [Ost92]. issues [Leb99]. Iterative [PTA^{+}92].

jellium [Bra93a]. John [Mer16, Mer93]. Jona [Kle92]. Josephson [Pac98]. jump [PK98]. justification [Nen91].

keV [Fra95]. Knot [Wu92, Wu93]. knots [Ati95]. Kolmogorov [SS94]. Kondo [TSU97]. Kramers [HTB90].

laboratory [Oka92]. Langmuir [KMD99]. large [Coh96]. large- [Coh96]. Laser [HW99, LSST99, Phi98, Slu99, CCNR91]. lasers [BS97a]. Lasinio [Kle92]. Lattice [RZ94, Man91, TSU97]. Lattice-gas [RZ94]. layers [Swa95]. leading [BBL96]. learning [WRB93]. Lecture [Chu98, CT98, Koh99b, Lau99, Phi98, Pop99, Sto99, Tsu99]. lensing [LBNZ98]. lepton [BCKK92]. lepton-hadron [BCKK92]. Leptonic [RB95]. Lie [KM91b]. life [Cah96]. lifetimes [PF95]. Light [GN99, Kol99, Ree94]. Light-meson [GN99]. limit [GKKR96]. limited [Cha93, MG98]. limits [CD94, Coh96]. Linear [Goe99]. liquid [Lee97, Lee98, ODB97]. liquids [DO99, Her91b, JNB96]. Local [ST92, MRR90]. localization [Rau92]. logarithms [BBL96]. Long [ODB97, AL97, Rob91]. Long-scale [ODB97]. long-term [AL97]. low [BK96a]. low-Q [BK96a]. low-X [BK96a].

MACHO [Sut99]. Macroscopic [Res94]. Magnetic [KBSE98, SGB^{+}95, YMS96, Gre98, Gre00, Nen91, OL97, PS96, Pou99, RT92, She94, She95, Tsu99]. magnetically [Gen95]. magnetization [Hir97]. magnetized [CCNR91]. Magnetoencephalography [HHI^{+}93]. magnetosphere [Win92]. mainly [Pei93]. manifestation [BU95]. Manipulating [CT98]. manipulation [Chu98, WPW99]. many [FS93, Sap98, Wir93]. many-body [FS93, Sap98, Wir93]. maps [Mei92]. masers [Ram90]. mass [JSSJ90]. material [YB95]. Materials [CD99, Her91a]. matrix [Bee97]. matter [Gol96, Koh99a, Koh99b, Sad99, Spr91, Wit99, dG92, dG99]. matter-wave [Koh99b]. mean [GKKR96]. mean-field [GKKR96]. measurement [BO96, MA91]. measurements [BK96b, CSB98, Ros92, Ros94]. mechanics [For92, Leb99, Omn92, WRB93, Wu92, Wu93]. mechanism [Fan92]. mechanisms [Bet90, KM94]. media [Isi92]. medical [Hen99]. medium [BKS95, HT99a]. melting [Das99]. meson [Ams98, GN99]. mesoscopic [Wei93]. mesoscopy [vRN99]. Metal [IFT98, Bra93a, BT99, dH93]. Metal-insulator [IFT98]. metals [BPS90, OT90, OL97, BPS91]. methods [Bra93b, Goe99, Sap98]. microcavities [KGJ^{+}99]. microlensing [Sut99]. Microscopic [Cep99, Hir97, ST92]. microscopy [SGB^{+}95, YB95, vRN99]. microwave [PW99]. millennium [TT99]. minimization [PTA^{+}92]. Mod [Gre00, JP90, KKL^{+}99a, Lee98, Mer16, Wei99a]. mode [CLvdB^{+}98, KGJ^{+}99, Swa95]. model [Bra93a, Fri91, Kle92, Rau92, TSU97, Cah96, GGS99, LLM92]. Modeling [JAP97]. models [BP92, BPW95, Coh96, Jen99, Pop99, RZ94, Sah93, Seg91, SFJ98, dH93]. modern [HK91, Sta99]. Molecular [Bis90, HKL99, Her99, Sis93, AM98, DDLÖ94, JAP97, Mea92, PTA^{+}92]. Molecular-beam [Sis93]. molecule [BT99]. molecule-metal [BT99]. molecules [BTS98, PR92, RKMG92]. moment [BS91, BS92]. monoclinic [Mer92a, Mer92b, Mer92c, Mer94]. monolayer [KBSE98]. monolayers [KMD99]. monopoles [LBNZ98]. Moon [Gut98]. most [Cro99]. motion [Bis90, Lis99, PSdH97, VM95]. motions [LR97]. motors [JAP97]. Mott [BK94]. Multichannel [AGLK96]. multiphase [RZ94]. Multiple [vRN99]. Multiresolution [Ari99]. muon [HK99]. muonium [SP98]. muons [SP98].

Nambu [Kle92]. nanokelvin [OL97]. nanosecond [Blo99]. nanostructures [AM98, Jen99, SB99]. narrow [MRR90]. narrow-band [MRR90]. national [Dre99]. Natural [BN98, Seg91]. nature [OC99, Sad99]. near [Vav94]. near-field [Vav94]. nearly [KM89, KM91a]. negative [BC94, OL97]. negative-ion [BC94]. networks [Hop99, NYO94]. neural [Hop99]. neutral [Chu98, DCCL91, Pau90, Phi98]. Neutrino [Wol99, BP92, CSB98, DSS90, Fra95, Kir99, Rei96, Rob92]. neutron [Bro95, Pet92, Shu95, Wir93]. neutrons [Gol96]. Newton [LBNZ98]. Nobel [Chu98, CT98, Koh99b, Lau99, Phi98, Pop99, Sto99, Tsu99]. noble [WH97]. noble-gas [WH97]. noise [HK96, Wei93]. non [DCCL91]. non-neutral [DCCL91]. nonclassical [Kol99]. nondemolition [BK96b]. nonequilibrium [GL99]. noninvasive [HHI^{+}93]. Nonlinear [AM98, Egg97, KGJ^{+}99, Mee96, Rob97]. nonneutral [DO99]. nonrelativistic [FS93]. nonretarded [MRR90]. nonuniform [BKS95]. normal [KGJ^{+}99]. normal-mode-coupling [KGJ^{+}99]. notes [Spr91]. Novel [CCDR^{+}93]. Nuclear [Bet99a, BS97b, Fow99, HS99, Ich93, JSH^{+}98, OL97, Ost92, PS96, Rii94, Til99, BT98, Gre98, Gre00, Iac93, KM91b, Pou99, Wir93]. Nucleation [SO95]. nuclei [BPP93, BN96, DDLÖ94, JSSJ90, JSH^{+}98, WH97]. Nucleon [LLL93, CS98]. Nucleon-nucleon [LLL93]. nucleosynthesis [Sal99, ST98]. nucleus [MBW99]. number [DSS90]. numerical [NYO94]. NUT [LBNZ98].

objects [AL97]. observation [Ken91]. observed [ABST93]. Observing [VM95]. odd [JSSJ90, JSH^{+}98]. odd- [JSSJ90]. odd-odd [JSH^{+}98]. oldest [Gut98]. one [KS96, Man99, TSU97]. one-dimensional [TSU97]. one-photon [Man99]. open [CLvdB^{+}98, Fre90, Fre91]. optical [HW96, KBSE98, WH97]. optically [Mee96]. optics [AM98, Dav96, HW99, KGJ^{+}99, PK98]. Optogalvanic [BBS90]. orbits [BBF^{+}94]. ordering [OL97, SB99]. orientationally [LBM94]. origin [Bru90, Ree94, BU95]. origins [Fit99]. orthorhombic [Mer92a, Mer92b, Mer92c, Mer94]. oscillator [BO96]. oscillatory [Ram90]. other [KG99]. outer [Joh96]. overview [LLK^{+}97]. oxides [KBSE98, Man91].

pair [Kle99a]. pairing [SR95, Van95]. paper [JP90]. paradigm [Swa95]. paramagnetic [SR95]. parameters [Cah96, HKL99]. Parity [MBW99]. particle [Deh90, Fit99, GGS99, Pai99, PSdH97, Ram91, Rei96]. particles [CCF^{+}99, Chu98, Cro99, DROM92, Pau90]. Path [Cep95]. Pattern [CH93, GL99, KM94]. patterns [BN98]. Pedagogic [Spr91]. Penning [Sis93]. Percolation [Isi92, Sah93]. periodic [BBF^{+}94, Lif97]. personal [Rob92]. perspective [AM98]. perturbative [SSC^{+}95, SS96]. Phase [KS96, MO96, Van95, DL93, GL91, KMD99, LBM94, Mea92, Res94, RZ94, SGCS97, TSU97]. Phase-sensitive [Van95]. phases [Lee97, Lee98]. phenomena [Fan92, LBM94, Sah93, Seg91, SR95, Sta99]. Phenomenological [SU91]. Phonon [Qui97, BHK98]. phonons [BHK98]. Photodissociation [HT99a]. photon [Man99, RB98]. photonics [HK96]. photons [CT98]. Phys [Gre00, JP90, KKL^{+}99a, Lee98, Mer16, Wei99a]. Physical [DD95, ABST93, Bro95, For95, NYO94]. physicists [PW97]. Physics [BoP96, BL99, Dre99, FWA99, Hen99, MR99, PS93, AC99, Ati95, BCKK92, Bet99a, BGJ90, Bra93a, Cep99, CD99, DDG^{+}99, Fis98, Fit99, GGS99, GL99, HS99, Her99, Iac93, KM91b, Kle99b, Koh99a, LSST99, LLM92, Pai99, She94, She95, Wol99, Zei99, dH93, BoP96]. pillars [Sta99]. Planetary [Ost93]. plasma [BS97b]. plasmas [CCNR91, DO99, Gen95, Ich93, Mee96, OC99, Oka92]. ploughshares [For95]. Point [dVvCL98]. Poisson [Bav91]. Poissonian [Dav96]. polaritons [BHK98]. polarization [Res94]. polaron [GL91]. polaronic [GL91]. pole [MMRT99]. poltergeist [Rei96]. Polymer [Wit98, DL93]. Pomeranchuk [Ric97]. population [BTS98]. positive [OL97]. positrons [PN94]. potential [LLK^{+}97]. power [Fow99]. Precision [CSB98, MMRT99, LLM92, ST98]. pressure [kMH94]. Primary [BU95]. principles [Mei92]. probes [BHK98]. problem [BP92, Gut98, LR97, OT90, Sap98]. problems [DROM92]. processes [Dav96, DD95, NF95, SS94]. production [Kle99a, RKMG92]. Progress [BBF^{+}94, BO96, SM94, WIP^{+}97]. project [Rob92, Sut99]. propagation [BKS95]. Properties [Car90, Bav91, DCCL91, GL91, Her91a, KBSE98, NYO94]. Proton [Ams98, Ken91, RKMG92]. Proton-antiproton [Ams98]. pulsar [Hul94]. pulsars [Tay94]. pumping [WH97].

Q [BK96a]. QCD [SS98, Shu93, SSC^{+}95, SS96]. quantal [BU95]. quantization [Lau99]. quantized [NF95]. Quantum [Ati95, BT98, Bet99b, BK96b, CD94, EJ96, For92, HK96, LSST99, Man99, Pop99, Ram91, SP98, Wil99, Bee97, BTS98, CCDR^{+}93, Coh96, Dav96, Fre90, Fre91, Gre98, Gre00, HW99, Huc95, Kle92, KKL^{+}99a, KKL^{+}99b, MG98, MO96, Omn92, PK98, SGCS97, STG99, Sto99, WPW99, Zei99]. quantum-jump [PK98]. quark [CF97, Cve99, Fri91, Ros92, Ros94]. Quasicrystals [GK93b, Mer92a, Mer92b, Mer92c, Mer94, Qui97, RMRW91, GK93a]. Quasinormal [CLvdB^{+}98]. Quasinormal-mode [CLvdB^{+}98]. quasiperiodic [Lif97].

R [Her91a]. radar [For95, Ost93, Pou99]. radiated [Vav94]. radiation [Cha93, PW99, Wei99a, Wei99b]. radiative [Mee96]. radio [SFJ98]. Random [Bee97, Eva93, Isi92, VG90]. Random-matrix [Bee97]. random-site [VG90]. Rare [RW93]. rate [HTB90]. rates [BS97b, CD94]. ray [BS97a, CT94, HSFP99, OT90]. rays [Cro99]. Reaction [HTB90, BS97b, MR99]. Reaction-rate [HTB90]. reactions [Mar93, MG98]. reactors [She94, Til99, She95]. realizations [KM91b, NYO94]. Recoilless [BS97a]. Reconstructing [LLK^{+}97]. refined [SS94]. reflection [BN96]. reflections [CT94]. region [BK96a]. regions [HT99a, JSSJ90]. relations [PF95]. relativistic [Sap98, Tay94]. Relativity [Sha99]. relaxation [YMS96]. Renormalization [Fis98, Sha94, Sta99]. Renormalization-group [Sha94]. reparametrization [MMTB96]. report [Sut99]. research [For95]. resistivities [BPS90, BPS91]. resolution [RWF97]. resonance [GHJM98, Gre98, Gre00, PS96, Pou99, SGB^{+}95]. Resonances [PR92, BC94, PF95]. response [Deg99]. rest [Deh90]. Results [Haw98, Kir99, SFJ98]. retina [RB98]. Rev [Gre00, JP90, KKL^{+}99a, Lee98, Mer16, Wei99a]. Review [BoP96, Leb99]. Riemannian [Rup95, Rup96]. rigorous [Nen91]. RMP [Pin92]. rocks [Sah93]. rod [RB98]. role [Gol96, Pac98]. rotation [LBM94]. rotational [Bis90]. rotations [LR97]. rotor [Rau92]. route [BK96b]. rule [WRB93]. Rydberg [AGLK96, Bra93b, CG99].

S187 [Wei99a]. sand [Kad99]. sandpile [Nag92]. scale [EPV95, ODB97]. Scaling [Huc95, NYO94, Sta99, Goe99, Ken91]. Scanning [YB95]. scatterers [dVvCL98]. scattering [BCKK92, BPP93, Fri91, FKT91, Ken91, LLL93, PR92, Shu95, Tay91, vRN99]. science [Blo99, BL93]. search [Das99, LLM92, WW93]. Searches [CCF^{+}99]. sections [A^{+}98, LLL93, SS96]. security [Dre99]. selective [Leb99]. self [Bra93a, Her91b]. self-consistent [Bra93a]. self-trapping [Her91b]. semicardinal [Ari99]. semiclassical [Bra93a, Bra93b]. semiconductor [AM98, KGJ^{+}99, SM90]. semiconductors [Van92]. semileptonic [RB95]. sensitive [Van95]. separated [Ram90]. separation [LR97, RZ94]. sequential [Eva93]. Shock [BKS95, MA91]. Shock-wave [BKS95, MA91]. Shor [EJ96]. short [Kle99b]. similarity [SS94]. simple [Bra93a, Jen99, OT90, OL97, dH93]. simulated [Sah93]. simulations [Cep99, NYO94]. Single [RB98, Kas92, Ram91]. single-electron [Kas92]. single-particle [Ram91]. Single-photon [RB98]. site [VG90]. Slow [Bro95]. Small [BCKK92]. Small- [BCKK92]. Soft [dG92, OT90, Wit99]. soft-X-ray [OT90]. Solar [A^{+}98, BPW95, Kir99, BP92, Bru90, Joh96, Lis99]. solid [kMH94, PN94]. solids [JNB96, MBB^{+}92, Pei93, PN94, Ram91, ST92, SP98, VM95]. Solitons [HW96, KM89, KM91a]. solutions [Wit98]. some [Spr91, ZFG90]. space [Kro97, LBNZ98, Mer92a, Mer92b, Mer92c, Mer94, Oka92, RMRW91]. spatial [Kol99]. species [DSS90]. spectra [LBNZ98]. spectroscopy [Ams98, AGLK96, BBS90, Bro95, CMD94a, GN99, HW99, CMD94b]. spectrum [Oka92]. spherical [Vav94]. Spin [FAG^{+}94, WH97, For92, Grü94, Man91, Ost92, Wei93]. spin- [Man91]. spin-density [Grü94]. Spin-density-wave [FAG^{+}94]. Spin-exchange [WH97]. Spontaneous [SB99]. Sputtering [Joh96]. square [Man91]. SR [Ama97]. stability [DCCL91, Spr91]. Standard [BP92, Cah96, GGS99, LLM92]. stars [Kro97, Pet92, Spr91, WIP^{+}97, Wir93]. state [LCD^{+}87, LCD^{+}95, TSU97, Van95]. states [BTS98, DO99, JSSJ90, Rii94, ZFG90]. static [Bis90]. Statistical [Leb99, Fis98, Isi92, WRB93, Wu92, Wu93]. statistics [For92]. Stellar [Sal99]. Stochastic [GHJM98, MMTB96, SS94]. String [Pol96, SS99]. strong [Rob97]. strongly [GKKR96]. Structure [CS98, KMD99, Ari99, Goe99, JSH^{+}98, Koh99b, SM90]. structures [KM94]. studied [Ama97]. studies [BPP93, Gol96, HHI^{+}93, Sis93]. study [DROM92]. Studying [TBW99]. Sub [Dav96]. Sub-Poissonian [Dav96]. subatomic [BGJ90, Deh90]. Summary [BoP96]. Sun [Gut98]. superconducting [Gre98, Gre00, KKL^{+}99a, KKL^{+}99b]. Superconductivity [Gun97, MRR90, ST99, RT92, SU91]. superconductors [BFG^{+}94, Car90, Dag94, PS96, SR95, Van95, YMS96]. superfluid [NF95, Pac98, SO95]. Superfluidity [Leg99, Osh97]. superlattice [SM90]. Supernova [Bet90]. supersymmetric [CCF^{+}99]. supersymmetries [Iac93]. supersymmetry [SS99]. Suppression [Kle99a]. Surface [DL93, SO95, BT99, Egg97]. surfaces [PR92, PN94, SB99]. survey [OT90]. Swords [For95]. symmetries [Iac93]. Symmetry [Kro97, BU95, Lif97, Van95]. Symplectic [Mei92]. Synthesis [WIP^{+}97]. system [Bru90, Joh96, LCD^{+}87, Lis99, LCD^{+}95]. systems [ABST93, Ama97, CS98, CLvdB^{+}98, DL93, DDLÖ94, Fre90, Fre91, GKKR96, GL91, Gil98, KM89, Mea92, MRR90, PSdH97, VG90, KM91a].

tails [Van92]. technique [Ama97, For95]. techniques [PTA^{+}92]. technological [Her91a]. technology [Slu99]. telescopes [RWF97]. temperature [BPS90, BPS91, BFG^{+}94, Dag94, Gen95, KKL^{+}99b, OC99, SR95, Van95, YMS96, KKL^{+}99a]. temperature-dependent [BPS90, BPS91]. temperatures [OL97]. tentative [dG99]. term [AL97]. tests [MMRT99, Van95]. Tevatron [CCF^{+}99]. Their [Gol96]. theorems [Mer93, Mer16]. Theoretical [BT99, Kad99, Pai99, RT92, Sap98, DDLÖ94]. Theories [Bru90]. Theory [CCNR91, DGPS99, Lif97, OT90, PN94, SM90, Van92, ZFG90, Bee97, Bet99b, Bra93b, Cep95, Cra91, Fis98, For92, FS93, GKKR96, GK99, HHI^{+}93, HTB90, Huc95, MG98, Ram91, Rup95, SS99, SU91, Spr91, WBZJ99, Wil99, Wir93, Wu92, Wu93, Mar93, Rup96]. thermal [DO99]. thermodynamic [Rup95, Rup96]. thin [Mee96, ODB97]. thirties [Pei93]. Thomas [Spr91]. Three [Sta99, Gut98]. three-body [Gut98]. Time [DDLÖ94, LM94, PF95]. Time-dependent [DDLÖ94]. time-energy [PF95]. tokamak [Haw98]. tools [BK96b]. Top [Cve99, CF97, Ros92, Ros94]. Top-quark [Cve99]. topography [Isi92]. Topological [Gil98]. topology [Ati95]. total [LLL93, PTA^{+}92, SS96]. total-energy [PTA^{+}92]. touch [BR99]. toys [BK96b]. transfer [BTS98, Mar93]. transformation [EPV95]. transistor [Kas92, RHH99]. transition [BK94, KKL^{+}99a, KKL^{+}99b, SO95]. transitions [CA99, DL93, GL91, IFT98, KMD99, LBM94, kMH94, MO96, RZ94, SGCS97]. Translation [LBM94]. Translation-rotation [LBM94]. transmission [IL99]. transport [BH98, Bee97, Hor99, Isi92, KBSE98, Mei92, Ram91]. transverse [TBW99]. Trapped [DO99, DGPS99]. trapping [Her91b, Phi98, WPW99]. traps [Pau90]. treatments [DDLÖ94]. triclinic [Mer92a, Mer92b, Mer92c, Mer94]. tritium [Haw98]. tunneling [BT98, LM94]. turbulence [BH98, Rob97, Sre99, SS94]. twentieth [Koh99a, Kle99b]. two [Leb99, LCD^{+}87, LCD^{+}95, Man99, Mer93, Mer16, Tsu99]. two-dimensional [Tsu99]. two-photon [Man99]. two-state [LCD^{+}87, LCD^{+}95].

U.S. [Dre99]. Ultracold [Gol96, WBZJ99]. Ultrahigh [kMH94]. Ultrahigh-pressure [kMH94]. uncertainty [PF95]. unconventional [SU91]. understanding [Osh97]. unification [SS99]. universality [Sta99]. universe [AL97, Cro99]. Unusual [SR95]. upon [Kad99]. uses [MG98]. using [TBW99].

vacuum [KG99, Shu93]. values [HK99]. variables [Mer93, Mer16]. variational [Mei92]. variations [Wir93]. very [RT92, Rob91]. very-long-baseline [Rob91]. via [Wei93]. vibrational [Bis90]. view [dG99]. viewed [IL99]. violation [MBW99, WW93]. viscosity [MA91]. vision [Jul91]. Vlasov [Bav91]. Vortices [BFG^{+}94, NF95].

War [For95]. wave [BKS95, FAG^{+}94, Koh99b, MA91, Rob97, SR95, Vav94]. wavelet [Ari99]. wavelets [BN98]. waves [CLvdB^{+}98, Grü94, Hor99, JP89, JP90, MR99, dVvCL98, vRN99]. Weak [BBL96, BO96]. welding [DD95]. wetting [CCDR^{+}93]. widths [SS96]. without [BP92]. work [Pei93]. working [HHI^{+}93]. world [Bro95, CG99, For95].

X [BK96a, CT94, OT90]. X-ray [CT94].

years [EH99, HTB90, Tay91, WIP^{+}97].

References

Adelberger:1998:SFC

[A^{+}98] Eric G. Adelberger et al. Solar fusion cross sections. Reviews of Modern Physics, 70(4):1265–1291, October 1998. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.70.1265;   http://rmp.aps.org/abstract/RMP/
v70/i4/p1265_1.

Abarbanel:1993:AOC

[ABST93] Henry D. I. Abarbanel, Reggie Brown, John J. Sidorowich, and Lev Sh. Tsimring. The analysis of observed chaotic data in physical systems. Reviews of Modern Physics, 65(4): 1331–1392, October 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.65.1331;   http://
rmp.aps.org/abstract/RMP/v65/i4/p1331_1.

Abramowicz:1999:HCP

[AC99] Halina Abramowicz and Allen C. Caldwell. HERA collider physics. Reviews of Modern Physics, 71(5):1275–1409, October 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.1275;   http://rmp.aps.org/abstract/RMP/
v71/i5/p1275_1.

Aymar:1996:MRS

[AGLK96] Mireille Aymar, Chris H. Greene, and Eliane Luc-Koenig. Multichannel Rydberg spectroscopy of complex atoms. Reviews of Modern Physics, 68(4):1015–1123, October 1996. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.68.1015;   http://rmp.aps.org/abstract/RMP/
v68/i4/p1015_1.

Adams:1997:DUL

[AL97] Fred C. Adams and Gregory Laughlin. A dying universe: the long-term fate and evolution of astrophysical objects. Reviews of Modern Physics, 69(2):337–372, April 1997. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.69.337;
 http://rmp.aps.org/abstract/RMP/v69/i2/p337_1.

Axt:1998:NOS

[AM98] V. M. Axt and S. Mukamel. Nonlinear optics of semiconductor and molecular nanostructures; a common perspective. Reviews of Modern Physics, 70(1):145–174, January 1998. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.70.145;
 http://rmp.aps.org/abstract/RMP/v70/i1/p145_1.

Amato:1997:HFS

[Ama97] A. Amato. Heavy-fermion systems studied by μSR technique. Reviews of Modern Physics, 69(4):1119–1180, October 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.69.1119;   http://rmp.aps.org/abstract/RMP/
v69/i4/p1119_1.

Amsler:1998:PAA

[Ams98] Claude Amsler. Proton-antiproton annihilation and meson spectroscopy with the Crystal Barrel. Reviews of Modern Physics, 70 (4):1293–1339, October 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.70.1293;   http://
rmp.aps.org/abstract/RMP/v70/i4/p1293_1.

Anselmino:1993:D

[APE^{+}93] Mauro Anselmino, Enrico Predazzi, Svante Ekelin, Sverker Fredriksson, and D. B. Lichtenberg. Diquarks. Reviews of Modern Physics, 65(4):1199–1233, October 1993. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.65.1199;   http://rmp.aps.org/abstract/RMP/
v65/i4/p1199_1.

Arias:1999:MAE

[Ari99] T. A. Arias. Multiresolution analysis of electronic structure: semicardinal and wavelet bases. Reviews of Modern Physics, 71(1):267–311, January 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.71.267;   http://
rmp.aps.org/abstract/RMP/v71/i1/p267_1.

Atiyah:1995:QPT

[Ati95] Michael Atiyah. Quantum physics and the topology of knots. Reviews of Modern Physics, 67(4):977–981, October 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
67.977;   http://rmp.aps.org/abstract/RMP/v67/i4/p977_
1.

Bavaud:1991:EPV

[Bav91] François Bavaud. Equilibrium properties of the Vlasov functional: The generalized Poisson–Boltzmann–Emden equation. Reviews of Modern Physics, 63(1):129–149, January 1991. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.63.129;
 http://rmp.aps.org/abstract/RMP/v63/i1/p129_1.

Badii:1994:PAE

[BBF^{+}94] R. Badii, E. Brun, M. Finardi, L. Flepp, R. Holzner, J. Parisi, C. Reyl, and J. Simonet. Progress in the analysis of experimental chaos through periodic orbits. Reviews of Modern Physics, 66 (4):1389–1415, October 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.66.1389;   http://
rmp.aps.org/abstract/RMP/v66/i4/p1389_1.

Buchalla:1996:WDB

[BBL96] Gerhard Buchalla, Andrzej J. Buras, and Markus E. Lautenbacher. Weak decays beyond leading logarithms. Reviews of Modern Physics, 68(4):1125–1244, October 1996. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/

RevModPhys.68.1125;   http://rmp.aps.org/abstract/RMP/
v68/i4/p1125_1.

Barbieri:1990:OS

[BBS90] Beniamino Barbieri, Nicolò Beverini, and Antonio Sasso. Optogalvanic spectroscopy. Reviews of Modern Physics, 62(3):603–644, July 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.62.603;   http://rmp.aps.org/abstract/
RMP/v62/i3/p603_1.

Buckman:1994:ANI

[BC94] Stephen J. Buckman and Charles W. Clark. Atomic negativeion resonances. Reviews of Modern Physics, 66(2):539–655, April 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.66.539;   http://rmp.aps.org/abstract/
RMP/v66/i2/p539_1.

Badelek:1992:SPD

[BCKK92] B. Badelek, K. Charchula, M. Krawczyk, and J. Kwieciński. Smallx physics in deep-inelastic lepton-hadron scattering. Reviews of Modern Physics, 64(4):927–960, October 1992. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.64.927;
 http://rmp.aps.org/abstract/RMP/v64/i4/p927_1.

Beenakker:1997:RMT

[Bee97] C. W. J. Beenakker. Random-matrix theory of quantum transport. Reviews of Modern Physics, 69(3):731–808, July 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
69.731;   http://rmp.aps.org/abstract/RMP/v69/i3/p731_
1.

Bethe:1990:SM

[Bet90] H. A. Bethe. Supernova mechanisms. Reviews of Modern Physics, 62(4):801–866, October 1990. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.62.801;
 http://rmp.aps.org/abstract/RMP/v62/i4/p801_1.

Bethe:1999:NP

[Bet99a] Hans A. Bethe. Nuclear physics. Reviews of Modern Physics, 71(2):S6–S15, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S6;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS6_1.

Bethe:1999:QT

[Bet99b] Hans A. Bethe. Quantum theory. Reviews of Modern Physics, 71(2):S1–S5, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S1;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS1_1.

Blatter:1994:VHT

[BFG^{+}94] G. Blatter, M. V. Feigel’man, V. B. Geshkenbein, A. I. Larkin, and V. M. Vinokur. Vortices in high-temperature superconductors. Reviews of Modern Physics, 66(4):1125–1388, October 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.66.1125;   http://rmp.aps.org/abstract/RMP/
v66/i4/p1125_1.

Boal:1990:IIS

[BGJ90] David H. Boal, Claus-Konrad Gelbke, and Byron K. Jennings. Intensity interferometry in subatomic physics. Reviews of Modern Physics, 62(3):553–602, July 1990. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.62.553;
 http://rmp.aps.org/abstract/RMP/v62/i3/p553_1.

Balbus:1998:ITE

[BH98] Steven A. Balbus and John F. Hawley. Instability, turbulence, and enhanced transport in accretion disks. Reviews of Modern Physics, 70(1):1–53, January 1998. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.70.1;
http://rmp.aps.org/abstract/RMP/v70/i1/p1_1.

Bakker:1998:CPP

[BHK98] H. J. Bakker, S. Hunsche, and H. Kurz. Coherent phonon polaritons as probes of anharmonic phonons in ferroelectrics. Reviews of Modern Physics, 70(2):523–536, April 1998. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.70.523;
 http://rmp.aps.org/abstract/RMP/v70/i2/p523_1.

Bishop:1990:MVR

[Bis90] David M. Bishop. Molecular vibrational and rotational motion in static and dynamic electric fields. Reviews of Modern Physics, 62(2):343–374, April 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.62.343;   http://
rmp.aps.org/abstract/RMP/v62/i2/p343_1.

Belitz:1994:AMT

[BK94] D. Belitz and T. R. Kirkpatrick. The Anderson–Mott transition. Reviews of Modern Physics, 66(2):261–380, April 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
66.261;   http://rmp.aps.org/abstract/RMP/v66/i2/p261_
1.

Badelek:1996:LQL

[BK96a] B. Badelek and J. Kwieciński. The low-Q^{2}, low-X region in electroproduction. Reviews of Modern Physics, 68(2):445–471, April 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.68.445;   http://rmp.aps.org/abstract/
RMP/v68/i2/p445_1.

Braginsky:1996:QNM

[BK96b] V. B. Braginsky and F. Ya. Khalili. Quantum nondemolition measurements: the route from toys to tools. Reviews of Modern Physics, 68(1):1–11, January 1996. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.68.1;
http://rmp.aps.org/abstract/RMP/v68/i1/p1_1.

Bisnovatyi-Kogan:1995:SWP

[BKS95] G. S. Bisnovatyi-Kogan and S. A. Silich. Shock-wave propagation in the nonuniform interstellar medium. Reviews of Modern Physics, 67(3):661–712, July 1995. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.67.661;
 http://rmp.aps.org/abstract/RMP/v67/i3/p661_1.

Burrows:1993:SBD

[BL93] Adam Burrows and James Liebert. The science of brown dwarfs. Reviews of Modern Physics, 65(2):301–336, April 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
65.301;   http://rmp.aps.org/abstract/RMP/v65/i2/p301_
1.

Brinkman:1999:PCI

[BL99] W. F. Brinkman and D. V. Lang. Physics and the communications industry. Reviews of Modern Physics, 71(2): S480–S488, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S480;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS480_1.

Bloembergen:1999:NFS

[Blo99] N. Bloembergen. From nanosecond to femtosecond science. Reviews of Modern Physics, 71(2):S283–S287, March 1999. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S283;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS283_1.

Butler:1996:IRA

[BN96] P. A. Butler and W. Nazarewicz. Intrinsic reflection asymmetry in atomic nuclei. Reviews of Modern Physics, 68(2):349–421, April 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.68.349;   http://rmp.aps.org/abstract/
RMP/v68/i2/p349_1.

Bowman:1998:NPW

[BN98] C. Bowman and A. C. Newell. Natural patterns and wavelets. Reviews of Modern Physics, 70(1):289–301, January 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
70.289;   http://rmp.aps.org/abstract/RMP/v70/i1/p289_

Bocko:1996:MWC

[BO96] Mark F. Bocko and Roberto Onofrio. On the measurement of a weak classical force coupled to a harmonic oscillator: experimental progress. Reviews of Modern Physics, 68(3):755–799, July 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.68.755;   http://rmp.aps.org/abstract/
RMP/v68/i3/p755_1.

Barnett:1996:PPS

[BoP96] R. M. Barnett and others (Particle Data Group). Particle physics summary: a digest of the 1996 Review of Particle Physics. Reviews of Modern Physics, 68(3):611–732, July 1996. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.68.611;
 http://rmp.aps.org/abstract/RMP/v68/i3/p611_1.

Bahcall:1992:SSM

[BP92] J. N. Bahcall and M. H. Pinsonneault. Standard solar models, with and without helium diffusion, and the solar neutrino problem. Reviews of Modern Physics, 64(4):885–926, October 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
64.885;   http://rmp.aps.org/abstract/RMP/v64/i4/p885_
1.

Benhar:1993:ESS

[BPP93] O. Benhar, V. R. Pandharipande, and Steven C. Pieper. Electronscattering studies of correlations in nuclei. Reviews of Modern Physics, 65(3):817–828, July 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.65.817;   http:/
/rmp.aps.org/abstract/RMP/v65/i3/p817_1.

Bass:1990:TDE

[BPS90] Jack Bass, William P. Pratt, Jr., and Peter A. Schroeder. The temperature-dependent electrical resistivities of the alkali metals. Reviews of Modern Physics, 62(3):645–744, July 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
62.645;   http://rmp.aps.org/abstract/RMP/v62/i3/p645_

1. See erratum [BPS91].

Bass:1991:ETD

[BPS91] Jack Bass, William P. Pratt, Jr., and Peter A. Schroeder. Erratum: The temperature-dependent electrical resistivities of the alkali metals. Reviews of Modern Physics, 63(1):213, January 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.63.213;   http://rmp.aps.org/abstract/
RMP/v63/i1/p213_1. See [BPS90].

Bahcall:1995:SMH

[BPW95] John N. Bahcall, M. H. Pinsonneault, and G. J. Wasserburg. Solar models with helium and heavy-element diffusion. Reviews of Modern Physics, 67(4):781–808, October 1995. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.67.781;
 http://rmp.aps.org/abstract/RMP/v67/i4/p781_1.

Binnig:1999:TA

[BR99] G. Binnig and H. Rohrer. In touch with atoms. Reviews of Modern Physics, 71(2):S324–S330, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S324;   http:
//rmp.aps.org/abstract/RMP/v71/i2/pS324_1.

Brack:1993:PSM

[Bra93a] Matthias Brack. The physics of simple metal clusters: selfconsistent jellium model and semiclassical approaches. Reviews of Modern Physics, 65(3):677–732, July 1993. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.65.677;
 http://rmp.aps.org/abstract/RMP/v65/i3/p677_1.

Braun:1993:DSM

[Bra93b] P. A. Braun. Discrete semiclassical methods in the theory of Rydberg atoms in external fields. Reviews of Modern Physics, 65(1):115–161, January 1993. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.65.115;
 http://rmp.aps.org/abstract/RMP/v65/i1/p115_1.

Brockhouse:1995:SNS

[Bro95] Bertram N. Brockhouse. Slow neutron spectroscopy and the grand atlas of the physical world. Reviews of Modern Physics, 67(4):735–751, October 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.67.735;   http://
rmp.aps.org/abstract/RMP/v67/i4/p735_1.

Brush:1990:TOS

[Bru90] Stephen G. Brush. Theories of the origin of the solar system 1956– 1985. Reviews of Modern Physics, 62(1):43–112, January 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.62.43;   http://rmp.aps.org/abstract/RMP/v62/
i1/p43_1.

Bernreuther:1991:EDM

[BS91] Werner Bernreuther and Mahiko Suzuki. The electric dipole moment of the electron. Reviews of Modern Physics, 63 (2):313–340, April 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.63.313;   http:/
/rmp.aps.org/abstract/RMP/v63/i2/p313_1.  See erratum

[BS92].

Bernreuther:1992:EED

[BS92] Werner Bernreuther and Mahiko Suzuki. Erratum: The electric dipole moment of the electron. Reviews of Modern Physics, 64(2):633, April 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.64.633;   http://
rmp.aps.org/abstract/RMP/v64/i2/p633_1. See [BS91].

Baldwin:1997:RGR

[BS97a] George C. Baldwin and Johndale C. Solem. Recoilless gammaray lasers. Reviews of Modern Physics, 69(4):1085–1118, October 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.69.1085;   http://rmp.aps.org/abstract/
RMP/v69/i4/p1085_1.

Brown:1997:NRR

[BS97b] Lowell S. Brown and R. F. Sawyer. Nuclear reaction rates in a plasma. Reviews of Modern Physics, 69(2):411–436, April 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.69.411;   http://rmp.aps.org/abstract/
RMP/v69/i2/p411_1.

Balantekin:1998:QTN

[BT98] A. B. Balantekin and N. Takigawa. Quantum tunneling in nuclear fusion. Reviews of Modern Physics, 70(1):77–100, January 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.70.77;   http://rmp.aps.org/abstract/RMP/v70/
i1/p77_1.

Brivio:1999:AMM

[BT99] G. P. Brivio and M. I. Trioni. The adiabatic moleculemetal surface interaction: Theoretical approaches. Reviews of Modern Physics, 71(1):231–265, January 1999. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.71.231;
 http://rmp.aps.org/abstract/RMP/v71/i1/p231_1.

Bergmann:1998:CPT

[BTS98] K. Bergmann, H. Theuer, and B. W. Shore. Coherent population transfer among quantum states of atoms and molecules. Reviews of Modern Physics, 70(3):1003–1025, July 1998. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.70.1003;   http://rmp.aps.org/abstract/RMP/
v70/i3/p1003_1.

Bohr:1995:PMS

[BU95] Aage Bohr and Ole Ulfbeck. Primary manifestation of symmetry. Origin of quantal indeterminacy. Reviews of Modern Physics, 67(1):1–35, January 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.67.1;   http://rmp.aps.
org/abstract/RMP/v67/i1/p1_1.

Chakrabarti:1999:DTH

[CA99] Bikas K. Chakrabarti and Muktish Acharyya. Dynamic transitions and hysteresis. Reviews of Modern Physics, 71(3):847–859, April 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.71.847;   http://rmp.aps.org/abstract/
RMP/v71/i3/p847_1.

Cahn:1996:EAP

[Cah96] Robert N. Cahn. The eighteen arbitrary parameters of the Standard Model in your everyday life. Reviews of Modern Physics, 68(3):951–959, July 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.68.951;   http://
rmp.aps.org/abstract/RMP/v68/i3/p951_1.

Carbotte:1990:PBE

[Car90] J. P. Carbotte. Properties of boson-exchange superconductors. Reviews of Modern Physics, 62(4):1027–1157, October 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.62.1027;   http://rmp.aps.org/abstract/RMP/
v62/i4/p1027_1.

Cheng:1993:NWB

[CCDR^{+}93] E. Cheng, M. W. Cole, J. Dupont-Roc, W. F. Saam, and J. Treiner. Novel wetting behavior in quantum films. Reviews of Modern Physics, 65(2):557–567, April 1993. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.65.557;
 http://rmp.aps.org/abstract/RMP/v65/i2/p557_1.

Carena:1999:SSP

[CCF^{+}99] M. Carena, R. Culbertson, H. Frisch, S. Eno, and S. Mrenna. Searches for supersymmetric particles at the Tevatron collider. Reviews of Modern Physics, 71(4):937–981, July 1999. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.71.937;
 http://rmp.aps.org/abstract/RMP/v71/i4/p937_1.

Cohen:1991:TFE

[CCNR91] Bruce I. Cohen, Ronald H. Cohen, William McCay Nevins, and Thomas D. Rognlien. Theory of free-electron-laser heating and current drive in magnetized plasmas. Reviews of Modern Physics, 63(4):949–990, October 1991. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.63.949;
 http://rmp.aps.org/abstract/RMP/v63/i4/p949_1.

Caves:1994:QLB

[CD94] Carlton M. Caves and P. D. Drummond. Quantum limits on bosonic communication rates. Reviews of Modern Physics, 66 (2):481–537, April 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.66.481;   http://
rmp.aps.org/abstract/RMP/v66/i2/p481_1.

Chaudhari:1999:MP

[CD99] P. Chaudhari and M. S. Dresselhaus. Materials physics. Reviews of Modern Physics, 71(2):S331–S335, March 1999. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S331;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS331_1.

Ceperley:1995:PIT

[Cep95] D. M. Ceperley. Path integrals in the theory of condensed helium. Reviews of Modern Physics, 67(2):279–355, April 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
67.279;   http://rmp.aps.org/abstract/RMP/v67/i2/p279_
1.

Ceperley:1999:MSP

[Cep99] D. M. Ceperley. Microscopic simulations in physics. Reviews of Modern Physics, 71(2):S438–S443, March 1999. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S438;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS438_1.

Campagnari:1997:DTQ

[CF97] Claudio Campagnari and Melissa Franklin. The discovery of the top quark. Reviews of Modern Physics, 69(1):137–212, January 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.69.137;   http://rmp.aps.org/abstract/
RMP/v69/i1/p137_1.

Clark:1999:ARE

[CG99] William Clark and Chris H. Greene. Adventures of a Rydberg electron in an anisotropic world. Reviews of Modern Physics, 71(3):821–833, April 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.71.821;   http://
rmp.aps.org/abstract/RMP/v71/i3/p821_1.

Cross:1993:PFO

[CH93] M. C. Cross and P. C. Hohenberg. Pattern formation outside of equilibrium. Reviews of Modern Physics, 65(3):851–1112, July 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.65.851;   http://rmp.aps.org/abstract/
RMP/v65/i3/p851_1.

Charpak:1993:EII

[Cha93] Georges Charpak. Electronic imaging of ionizing radiation with limited avalanches in gases. Reviews of Modern Physics, 65(3):591–598, July 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.65.591;   http://
rmp.aps.org/abstract/RMP/v65/i3/p591_1.

Chu:1998:NLM

[Chu98] Steven Chu. Nobel Lecture: The manipulation of neutral particles. Reviews of Modern Physics, 70(3):685–706, July 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
70.685;   http://rmp.aps.org/abstract/RMP/v70/i3/p685_
1.

Carlson:1994:DEF

[CLS94] J. M. Carlson, J. S. Langer, and B. E. Shaw. Dynamics of earthquake faults. Reviews of Modern Physics, 66(2):657–670, April 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.66.657;   http://rmp.aps.org/abstract/
RMP/v66/i2/p657_1.

Ching:1998:QME

[CLvdB^{+}98] E. S. C. Ching, P. T. Leung, A. Maassen van den Brink, W. M. Suen, S. S. Tong, and K. Young. Quasinormal-mode expansion for waves in open systems. Reviews of Modern Physics, 70 (4):1545–1554, October 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.70.1545;   http://
rmp.aps.org/abstract/RMP/v70/i4/p1545_1.

Coplan:1994:S

[CMD94a] M. A. Coplan, J. H. Moore, and J. P. Doering. (e, 2e) spectroscopy. Reviews of Modern Physics, 66(3):985–1014, July 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
66.985;   http://rmp.aps.org/abstract/RMP/v66/i3/p985_

1. See erratum [CMD94b].

Coplan:1994:ES

[CMD94b] M. A. Coplan, J. H. Moore, and J. P. Doering. Erratum: (e, 2e) spectroscopy. Reviews of Modern Physics, 66 (4):1517, October 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.66.1517;   http://
rmp.aps.org/abstract/RMP/v66/i4/p1517_1. See [CMD94a].

Cohen:1996:CLL

[Coh96] Thomas D. Cohen. Chiral and large-N_{c} limits of quantum chromodynamics and models of the baryon. Reviews of Modern Physics, 68(2):599–608, April 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.68.599;   http:/
/rmp.aps.org/abstract/RMP/v68/i2/p599_1.

Crawford:1991:IBT

[Cra91] John David Crawford. Introduction to bifurcation theory. Reviews of Modern Physics, 63(4):991–1037, October 1991. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.63.991;
 http://rmp.aps.org/abstract/RMP/v63/i4/p991_1.

Cronin:1999:CRM

[Cro99] James W. Cronin. Cosmic rays: the most energetic particles in the universe. Reviews of Modern Physics, 71(2): S165–S172, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S165;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS165_1.

Carlson:1998:SDF

[CS98] J. Carlson and R. Schiavilla. Structure and dynamics of fewnucleon systems. Reviews of Modern Physics, 70(3):743–841, July 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.70.743;   http://rmp.aps.org/abstract/
RMP/v70/i3/p743_1.

Conrad:1998:PMH

[CSB98] Janet M. Conrad, Michael H. Shaevitz, and Tim Bolton. Precision measurements with high-energy neutrino beams. Reviews of Modern Physics, 70(4):1341–1392, October 1998. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.70.1341;   http://rmp.aps.org/abstract/RMP/
v70/i4/p1341_1.

Carra:1994:AXR

[CT94] Paolo Carra and B. T. Thole. Anisotropic X-ray anomalous diffraction and forbidden reflections. Reviews of Modern Physics, 66 (4):1509–1515, October 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.66.1509;   http://
rmp.aps.org/abstract/RMP/v66/i4/p1509_1.

Cohen-Tannoudji:1998:NLM

[CT98] Claude N. Cohen-Tannoudji. Nobel Lecture: Manipulating atoms with photons. Reviews of Modern Physics, 70(3):707–719, July 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.70.707;   http://rmp.aps.org/abstract/
RMP/v70/i3/p707_1.

Curl:1997:DFE

[Cur97] Robert F. Curl. Dawn of the fullerenes: experiment and conjecture. Reviews of Modern Physics, 69(3):691–702, July 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.69.691;   http://rmp.aps.org/abstract/
RMP/v69/i3/p691_1.

Cvetic:1999:TQC

[Cve99] G. Cvetic. Top-quark condensation. Reviews of Modern Physics, 71(3):513–574, April 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.71.513;   http://
rmp.aps.org/abstract/RMP/v71/i3/p513_1.

Dagotto:1994:CEH

[Dag94] Elbio Dagotto. Correlated electrons in high-temperature superconductors. Reviews of Modern Physics, 66(3):763–840, July 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.66.763;   http://rmp.aps.org/abstract/
RMP/v66/i3/p763_1.

Dash:1999:HSC

[Das99] J. G. Dash. History of the search for continuous melting. Reviews of Modern Physics, 71(5):1737–1743, October 1999. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.1737;   http://rmp.aps.org/abstract/RMP/
v71/i5/p1737_1.

Davidovich:1996:SPP

[Dav96] Luiz Davidovich. Sub-Poissonian processes in quantum optics. Reviews of Modern Physics, 68(1):127–173, January 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
68.127;   http://rmp.aps.org/abstract/RMP/v68/i1/p127_
1.

Davidson:1991:ESP

[DCCL91] Ronald C. Davidson, Hei-Wai Chan, Chiping Chen, and Steven Lund. Equilibrium and stability properties of intense non-neutral electron flow. Reviews of Modern Physics, 63(2):341–374, April 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.63.341;   http://rmp.aps.org/abstract/
RMP/v63/i2/p341_1.

DebRoy:1995:PPF

[DD95] T. DebRoy and S. A. David. Physical processes in fusion welding. Reviews of Modern Physics, 67(1):85–112, January 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
67.85;   http://rmp.aps.org/abstract/RMP/v67/i1/p85_1.

Datz:1999:AP

[DDG^{+}99] Sheldon Datz, G. W. F. Drake, T. F. Gallagher, H. Kleinpoppen, and G. zu Putlitz. Atomic physics. Reviews of Modern Physics, 71(2):S223–S241, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S223;   http:
//rmp.aps.org/abstract/RMP/v71/i2/pS223_1.

Deumens:1994:TDT

[DDLÖ94] E. Deumens, A. Diz, R. Longo, and Y. Öhrn. Time-dependent theoretical treatments of the dynamics of electrons and nuclei in molecular systems. Reviews of Modern Physics, 66(3):917–983, July 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.66.917;   http://rmp.aps.org/abstract/
RMP/v66/i3/p917_1.

Degiorgi:1999:ERH

[Deg99] L. Degiorgi. The electrodynamic response of heavy-electron compounds. Reviews of Modern Physics, 71(3):687–734, April 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.71.687;   http://rmp.aps.org/abstract/
RMP/v71/i3/p687_1.

Dehmelt:1990:EIS

[Deh90] Hans Dehmelt. Experiments with an isolated subatomic particle at rest. Reviews of Modern Physics, 62(3):525–530, July 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.62.525;   http://rmp.aps.org/abstract/
RMP/v62/i3/p525_1.

deGennes:1992:SM

[dG92] P. G. de Gennes. Soft matter. Reviews of Modern Physics, 64(3):645–648, July 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.64.645;   http://
rmp.aps.org/abstract/RMP/v64/i3/p645_1.

deGennes:1999:GMT

[dG99] P. G. de Gennes. Granular matter: a tentative view. Reviews of Modern Physics, 71(2):S374–S382, March 1999. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S374;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS374_1.

Dalfovo:1999:TBE

[DGPS99] Franco Dalfovo, Stefano Giorgini, Lev P. Pitaevskii, and Sandro Stringari. Theory of Bose–Einstein condensation in trapped gases. Reviews of Modern Physics, 71(3):463–512, April 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
71.463;   http://rmp.aps.org/abstract/RMP/v71/i3/p463_
1.

deHeer:1993:PSM

[dH93] Walt A. de Heer. The physics of simple metal clusters: experimental aspects and simple models. Reviews of Modern Physics, 65(3):611–676, July 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.65.611;   http://
rmp.aps.org/abstract/RMP/v65/i3/p611_1.

DeBell:1993:SPT

[DL93] K. De’Bell and Turab Lookman. Surface phase transitions in polymer systems. Reviews of Modern Physics, 65(1): 87–113, January 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.65.87;   http://
rmp.aps.org/abstract/RMP/v65/i1/p87_1.

Dubin:1999:TNP

[DO99] Daniel H. E. Dubin and T. M. O’Neil. Trapped nonneutral plasmas, liquids, and crystals (the thermal equilibrium states). Reviews of Modern Physics, 71(1):87–172, January 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
71.87;   http://rmp.aps.org/abstract/RMP/v71/i1/p87_1.

Drell:1999:PUN

[Dre99] Sidney D. Drell. Physics and U.S. national security. Reviews of Modern Physics, 71(2):S460–S470, March 1999. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S460;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS460_1.

Darling:1992:FCP

[DROM92] T. W. Darling, F. Rossi, G. I. Opat, and G. F. Moorhead. The fall of charged particles under gravity: a study of experimental problems. Reviews of Modern Physics, 64(1):237–257, January 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.64.237;   http://rmp.aps.org/abstract/
RMP/v64/i1/p237_1.

Denegri:1990:NNS

[DSS90] D. Denegri, B. Sadoulet, and M. Spiro. The number of neutrino species. Reviews of Modern Physics, 62(1):1–42, January 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.aps.

org/doi/10.1103/RevModPhys.62.1;   http://rmp.aps.org/
abstract/RMP/v62/i1/p1_1.

deVries:1998:PSC

[dVvCL98] Pedro de Vries, David V. van Coevorden, and Ad Lagendijk. Point scatterers for classical waves. Reviews of Modern Physics, 70(2):447–466, April 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.70.447;   http://
rmp.aps.org/abstract/RMP/v70/i2/p447_1.

Eggers:1997:NDB

[Egg97] Jens Eggers. Nonlinear dynamics and breakup of free-surface flows. Reviews of Modern Physics, 69(3):865–930, July 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
69.865;   http://rmp.aps.org/abstract/RMP/v69/i3/p865_

Eades:1999:FYA

[EH99] J. Eades and F. J. Hartmann. Forty years of antiprotons. Reviews of Modern Physics, 71(1):373–419, January 1999. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.71.373;
 http://rmp.aps.org/abstract/RMP/v71/i1/p373_1.

Ekert:1996:QCS

[EJ96] Artur Ekert and Richard Jozsa. Quantum computation and Shor’s factoring algorithm. Reviews of Modern Physics, 68(3):733–753, July 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.68.733;   http://rmp.aps.org/abstract/
RMP/v68/i3/p733_1.

Erzan:1995:FST

[EPV95] A. Erzan, L. Pietronero, and A. Vespignani. The fixedscale transformation approach to fractal growth. Reviews of Modern Physics, 67(3):545–604, July 1995. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.67.545;
 http://rmp.aps.org/abstract/RMP/v67/i3/p545_1.

Evans:1993:RCS

[Eva93] J. W. Evans. Random and cooperative sequential adsorption. Reviews of Modern Physics, 65(4):1281–1329, October 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.65.1281;   http://rmp.aps.org/abstract/RMP/
v65/i4/p1281_1.

Fawcett:1994:SDW

[FAG^{+}94] E. Fawcett, H. L. Alberts, V. Yu. Galkin, D. R. Noakes, and J. V. Yakhmi. Spin-density-wave antiferromagnetism in chromium alloys. Reviews of Modern Physics, 66(1): 25–127, January 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.66.25;   http://
rmp.aps.org/abstract/RMP/v66/i1/p25_1.

Fano:1992:CMC

[Fan92] U. Fano. A common mechanism of collective phenomena. Reviews of Modern Physics, 64(1):313–319, January 1992. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.64.313;
 http://rmp.aps.org/abstract/RMP/v64/i1/p313_1.

Field:1999:A

[Fie99] George Field. Astrophysics. Reviews of Modern Physics, 71 (2):S33–S40, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.71.S33;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS33_1.

Fisher:1998:RGT

[Fis98] Michael E. Fisher. Renormalization group theory: Its basis and formulation in statistical physics. Reviews of Modern Physics, 70(2):653–681, April 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.70.653;   http://
rmp.aps.org/abstract/RMP/v70/i2/p653_1.

Fitch:1999:EPP

[Fit99] Val L. Fitch. Elementary particle physics: The origins. Reviews of Modern Physics, 71(2):S25–S32, March 1999. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.71.S25;
 http://rmp.aps.org/abstract/RMP/v71/i2/pS25_1.

Friedman:1991:DISb

[FKT91] Jerome I. Friedman, Henry W. Kendall, and Richard E. Taylor. Deep inelastic scattering: Acknowledgments. Reviews of Modern Physics, 63(3):629, July 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.63.629;   http:/
/rmp.aps.org/abstract/RMP/v63/i3/p629_1.

Forte:1992:QMF

[For92] Stefano Forte. Quantum mechanics and field theory with fractional spin and statistics. Reviews of Modern Physics, 64 (1):193–236, January 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.64.193;   http://
rmp.aps.org/abstract/RMP/v64/i1/p193_1.

Forman:1995:SPB

[For95] Paul Forman. “swords into ploughshares”: Breaking new ground with radar hardware and technique in physical research after World War II. Reviews of Modern Physics, 67(2):397–455, April 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.67.397;   http://rmp.aps.org/abstract/
RMP/v67/i2/p397_1.

Fowler:1999:NPF

[Fow99] T. Kenneth Fowler. Nuclear power — fusion. Reviews of Modern Physics, 71(2):S456–S459, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S456;   http:
//rmp.aps.org/abstract/RMP/v71/i2/pS456_1.

Franklin:1995:ADK

[Fra95] Allan Franklin. The appearance and disappearance of the 17keV neutrino. Reviews of Modern Physics, 67(2):457–490, April 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.67.457;   http://rmp.aps.org/abstract/
RMP/v67/i2/p457_1.

Frensley:1990:BCO

[Fre90] William R. Frensley. Boundary conditions for open quantum systems driven far from equilibrium. Reviews of Modern Physics, 62(3):745–791, July 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.62.745;   http:/
/rmp.aps.org/abstract/RMP/v62/i3/p745_1.  See erratum

[Fre91].

Frensley:1991:EBC

[Fre91] William R. Frensley. Erratum: Boundary conditions for open quantum systems driven far from equilibrium. Reviews of Modern Physics, 63(1):215, January 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.63.215;   http:/
/rmp.aps.org/abstract/RMP/v63/i1/p215_1. See [Fre90].

Friedman:1991:DISa

[Fri91] Jerome I. Friedman. Deep inelastic scattering: Comparisons with the quark model. Reviews of Modern Physics, 63(3):615–627, July 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.63.615;   http://rmp.aps.org/abstract/
RMP/v63/i3/p615_1.

Frohlich:1993:GIC

[FS93] Jürg Fröhlich and Urban M. Studer. Gauge invariance and current algebra in nonrelativistic many-body theory. Reviews of Modern Physics, 65(3):733–802, July 1993. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.65.733;
 http://rmp.aps.org/abstract/RMP/v65/i3/p733_1.

Frauenfelder:1999:BP

[FWA99] Hans Frauenfelder, Peter G. Wolynes, and Robert H. Austin. Biological physics. Reviews of Modern Physics, 71(2):S419– S430, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S419;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS419_1.

Gentle:1995:DMC

[Gen95] K. W. Gentle. Diagnostics for magnetically confined hightemperature plasmas. Reviews of Modern Physics, 67(4): 809–836, October 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.67.809;   http://
rmp.aps.org/abstract/RMP/v67/i4/p809_1.

Gaillard:1999:SMP

[GGS99] Mary K. Gaillard, Paul D. Grannis, and Frank J. Sciulli. The Standard Model of particle physics. Reviews of Modern Physics, 71(2):S96–S111, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.71.S96;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS96_1.

Gammaitoni:1998:SR

[GHJM98] Luca Gammaitoni, Peter Hänggi, Peter Jung, and Fabio Marchesoni. Stochastic resonance. Reviews of Modern Physics, 70 (1):223–287, January 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.70.223;   http://
rmp.aps.org/abstract/RMP/v70/i1/p223_1.

Gilmore:1998:TAC

[Gil98] Robert Gilmore. Topological analysis of chaotic dynamical systems. Reviews of Modern Physics, 70(4):1455–1529, October 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.70.1455;   http://rmp.aps.org/abstract/RMP/
v70/i4/p1455_1.

Goldman:1993:EQC

[GK93a] A. I. Goldman and K. F. Kelton. Erratum: Quasicrystals and crystalline approximants. Reviews of Modern Physics, 65(2):579, April 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.65.579;   http://rmp.aps.org/abstract/
RMP/v65/i2/p579_1. See [GK93b].

Goldman:1993:QCA

[GK93b] A. I. Goldman and R. F. Kelton. Quasicrystals and crystalline approximants. Reviews of Modern Physics, 65(1):213–230, January 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.65.213;   http://rmp.aps.org/abstract/

RMP/v65/i1/p213_1. See erratum [GK93a].

Giveon:1999:BDG

[GK99] Amit Giveon and David Kutasov. Brane dynamics and gauge theory. Reviews of Modern Physics, 71(4):983–1084, July 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.71.983;   http://rmp.aps.org/abstract/
RMP/v71/i4/p983_1.

Georges:1996:DMF

[GKKR96] Antoine Georges, Gabriel Kotliar, Werner Krauth, and Marcelo J. Rozenberg. Dynamical mean-field theory of strongly correlated fermion systems and the limit of infinite dimensions. Reviews of Modern Physics, 68(1):13–125, January 1996. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
68.13;   http://rmp.aps.org/abstract/RMP/v68/i1/p13_1.

Gerlach:1991:APP

[GL91] B. Gerlach and H. Löwen. Analytical properties of polaron systems or: Do polaronic phase transitions exist or not? Reviews of Modern Physics, 63(1):63–90, January 1991. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
63.63;   http://rmp.aps.org/abstract/RMP/v63/i1/p63_1.

Gollub:1999:PFN

[GL99] J. P. Gollub and J. S. Langer. Pattern formation in nonequilibrium physics. Reviews of Modern Physics, 71(2):S396– S403, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S396;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS396_1.

Godfrey:1999:LMS

[GN99] Stephen Godfrey and Jim Napolitano. Light-meson spectroscopy. Reviews of Modern Physics, 71(5):1411–1462, October 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.1411;   http://rmp.aps.org/abstract/RMP/
v71/i5/p1411_1.

Goedecker:1999:LSE

[Goe99] Stefan Goedecker. Linear scaling electronic structure methods. Reviews of Modern Physics, 71(4):1085–1123, July 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.1085;   http://rmp.aps.org/abstract/RMP/
v71/i4/p1085_1.

Golub:1996:UNT

[Gol96] R. Golub. Ultracold neutrons: Their role in studies of condensed matter. Reviews of Modern Physics, 68(2):329–347, April 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.68.329;   http://rmp.aps.org/abstract/
RMP/v68/i2/p329_1.

Greenberg:1998:ASQ

[Gre98] Ya. S. Greenberg. Application of superconducting quantum interference devices to nuclear magnetic resonance. Reviews of Modern Physics, 70(1):175–222, January 1998. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.70.175;
 http://rmp.aps.org/abstract/RMP/v70/i1/p175_1. See erra-

tum [Gre00].

Greenberg:2000:EAS

[Gre00] Ya. S. Greenberg. Erratum: Application of superconducting quantum interference devices to nuclear magnetic resonance [Rev. Mod. Phys. 70, 175 (1998)]. Reviews of Modern Physics, 72(1):329, January 2000. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.72.329;   http://rmp.aps.org/abstract/
RMP/v72/i1/p329_1. See [Gre98].

Gruner:1994:DSD

[Grü94] G. Grüner. The dynamics of spin-density waves. Reviews of Modern Physics, 66(1):1–24, January 1994. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.66.1;
http://rmp.aps.org/abstract/RMP/v66/i1/p1_1.

Gunnarsson:1997:SF

[Gun97] O. Gunnarsson. Superconductivity in fullerides. Reviews of Modern Physics, 69(2):575–606, April 1997. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.69.575;
 http://rmp.aps.org/abstract/RMP/v69/i2/p575_1.

Gutzwiller:1998:MES

[Gut98] Martin C. Gutzwiller. Moon–Earth–Sun: The oldest threebody problem. Reviews of Modern Physics, 70(2):589–639, April 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.70.589;   http://rmp.aps.org/abstract/
RMP/v70/i2/p589_1.

Hawryluk:1998:RDT

[Haw98] R. J. Hawryluk. Results from deuterium-tritium tokamak confinement experiments. Reviews of Modern Physics, 70 (2):537–587, April 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.70.537;   http://
rmp.aps.org/abstract/RMP/v70/i2/p537_1.

Hendee:1999:PAM

[Hen99] William R. Hendee. Physics and applications of medical imaging. Reviews of Modern Physics, 71(2):S444–S450, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S444;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS444_1.

Herbst:1991:RMI

[Her91a] J. F. Herbst. R_{2}Fe_{14}B materials: Intrinsic properties and technological aspects. Reviews of Modern Physics, 63(4):819–898, October 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.63.819;   http://rmp.aps.org/abstract/
RMP/v63/i4/p819_1.

Hernandez:1991:EST

[Her91b] John P. Hernandez. Electron self-trapping in liquids and dense gases. Reviews of Modern Physics, 63(3):675–697, July 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.63.675;   http://rmp.aps.org/abstract/
RMP/v63/i3/p675_1.

Herschbach:1999:CPM

[Her99] Dudley Herschbach. Chemical physics: Molecular clouds, clusters, and corrals. Reviews of Modern Physics, 71(2): S411–S418, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S411;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS411_1.

Hamalainen:1993:MTI

[HHI^{+}93] Matti Hämäläinen, Riitta Hari, Risto J. Ilmoniemi, Jukka Knuutila, and Olli V. Lounasmaa. Magnetoencephalography — theory, instrumentation, and applications to noninvasive studies of the working human brain. Reviews of Modern Physics, 65 (2):413–497, April 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.65.413;   http://
rmp.aps.org/abstract/RMP/v65/i2/p413_1.

Hirst:1997:MMC

[Hir97] L. L. Hirst. The microscopic magnetization: concept andapplication. Reviews of Modern Physics, 69(2):607–628, April 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.69.607;   http://rmp.aps.org/abstract/
RMP/v69/i2/p607_1.

Heifets:1991:CIM

[HK91] S. A. Heifets and S. A. Kheifets. Coupling impedance in modern accelerators. Reviews of Modern Physics, 63(3):631–673, July 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.63.631;   http://rmp.aps.org/abstract/
RMP/v63/i3/p631_1.

Henry:1996:QNP

[HK96] Charles H. Henry and Rudolf F. Kazarinov. Quantum noise in photonics. Reviews of Modern Physics, 68(3):801–853, July 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.68.801;   http://rmp.aps.org/abstract/
RMP/v68/i3/p801_1.

Hughes:1999:AVE

[HK99] V. W. Hughes and T. Kinoshita. Anomalous g values of the electron and muon. Reviews of Modern Physics, 71(2): S133–S139, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S133;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS133_1.

Harris:1999:MCC

[HKL99] A. B. Harris, Randall D. Kamien, and T. C. Lubensky. Molecular chirality and chiral parameters. Reviews of Modern Physics, 71(5):1745–1757, October 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.1745;   http://
rmp.aps.org/abstract/RMP/v71/i5/p1745_1.

Hoyle:1995:CAD

[HN95] F. Hoyle and J. V. Narlikar. Cosmology and action-at-a-distance electrodynamics. Reviews of Modern Physics, 67(1):113–155, January 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.67.113;   http://rmp.aps.org/abstract/
RMP/v67/i1/p113_1.

Hopfield:1999:BNN

[Hop99] J. J. Hopfield. Brain, neural networks, and computation. Reviews of Modern Physics, 71(2):S431–S437, March 1999. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S431;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS431_1.

Horton:1999:DWT

[Hor99] W. Horton. Drift waves and transport. Reviews of Modern Physics, 71(3):735–778, April 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.735;   http:/
/rmp.aps.org/abstract/RMP/v71/i3/p735_1.

Henley:1999:NPE

[HS99] E. M. Henley and J. P. Schiffer. Nuclear physics at the end of the century. Reviews of Modern Physics, 71(2):S205– S219, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S205;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS205_1.

Hoffman:1999:GRA

[HSFP99] C. M. Hoffman, C. Sinnis, P. Fleury, and M. Punch. Gammaray astronomy at high energies. Reviews of Modern Physics, 71(4):897–936, July 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.71.897;   http://
rmp.aps.org/abstract/RMP/v71/i4/p897_1.

Hollenbach:1999:PRI

[HT99a] D. J. Hollenbach and A. G. G. M. Tielens. Photodissociation regions in the interstellar medium of galaxies. Reviews of Modern Physics, 71(1):173–230, January 1999. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.71.173;
 http://rmp.aps.org/abstract/RMP/v71/i1/p173_1.

Horowitz:1999:BH

[HT99b] Gary T. Horowitz and Saul A. Teukolsky. Black holes. Reviews of Modern Physics, 71(2):S180–S186, March 1999. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S180;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS180_1.

Hanggi:1990:RRT

[HTB90] Peter Hänggi, Peter Talkner, and Michal Borkovec. Reaction-rate theory: fifty years after Kramers. Reviews of Modern Physics, 62(2):251–341, April 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.62.251;   http://
rmp.aps.org/abstract/RMP/v62/i2/p251_1.

Huckestein:1995:STI

[Huc95] Bodo Huckestein. Scaling theory of the integer quantum Hall effect. Reviews of Modern Physics, 67(2):357–396, April 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.67.357;   http://rmp.aps.org/abstract/
RMP/v67/i2/p357_1.

Hulse:1994:DBP

[Hul94] Russell A. Hulse. The discovery of the binary pulsar. Reviews of Modern Physics, 66(3):699–710, July 1994. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.66.699;
 http://rmp.aps.org/abstract/RMP/v66/i3/p699_1.

Haus:1996:SOC

[HW96] Hermann A. Haus and William S. Wong. Solitons in optical communications. Reviews of Modern Physics, 68(2):423–444, April 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.68.423;   http://rmp.aps.org/abstract/
RMP/v68/i2/p423_1.

Hansch:1999:LSQ

[HW99] T. W. Hänsch and H. Walther. Laser spectroscopy and quantum optics. Reviews of Modern Physics, 71(2):S242–S252, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S242;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS242_1.

Iachello:1993:DSS

[Iac93] F. Iachello. Dynamic symmetries and supersymmetries in nuclear physics. Reviews of Modern Physics, 65(2):569–576, April 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.65.569;   http://rmp.aps.org/abstract/
RMP/v65/i2/p569_1.

Ichimaru:1993:NFD

[Ich93] Setsuo Ichimaru. Nuclear fusion in dense plasmas. Reviews of Modern Physics, 65(2):255–299, April 1993. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.65.255;
 http://rmp.aps.org/abstract/RMP/v65/i2/p255_1.

Imada:1998:MIT

[IFT98] Masatoshi Imada, Atsushi Fujimori, and Yoshinori Tokura. Metal-insulator transitions. Reviews of Modern Physics, 70(4): 1039–1263, October 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.70.1039;   http://
rmp.aps.org/abstract/RMP/v70/i4/p1039_1.

Imry:1999:CVT

[IL99] Yoseph Imry and Rolf Landauer. Conductance viewed as transmission. Reviews of Modern Physics, 71(2):S306–S312, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S306;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS306_1.

Isichenko:1992:PST

[Isi92] M. B. Isichenko. Percolation, statistical topography, and transport in random media. Reviews of Modern Physics, 64(4): 961–1043, October 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.64.961;   http://
rmp.aps.org/abstract/RMP/v64/i4/p961_1.

Julicher:1997:MMM

[JAP97] Frank Jülicher, Armand Ajdari, and Jacques Prost. Modeling molecular motors. Reviews of Modern Physics, 69(4): 1269–1282, October 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.69.1269;   http://
rmp.aps.org/abstract/RMP/v69/i4/p1269_1.

Jensen:1999:GNC

[Jen99] Pablo Jensen. Growth of nanostructures by cluster deposition: Experiments and simple models. Reviews of Modern Physics, 71 (5):1695–1735, October 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.1695;   http://
rmp.aps.org/abstract/RMP/v71/i5/p1695_1.

Jaeger:1996:GSL

[JNB96] Heinrich M. Jaeger, Sidney R. Nagel, and Robert P. Behringer. Granular solids, liquids, and gases. Reviews of Modern Physics, 68(4):1259–1273, October 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.68.1259;   http://
rmp.aps.org/abstract/RMP/v68/i4/p1259_1.

Johnson:1996:SIO

[Joh96] R. E. Johnson. Sputtering of ices in the outer solar system. Reviews of Modern Physics, 68(1):305–312, January 1996. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.68.305;
 http://rmp.aps.org/abstract/RMP/v68/i1/p305_1.

Joseph:1989:HW

[JP89] D. D. Joseph and Luigi Preziosi. Heat waves. Reviews of Modern Physics, 61(1):41–73, January 1989. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.61.41;
 http://rmp.aps.org/abstract/RMP/v61/i1/p41_1. See ad-

dendum [JP90].

Joseph:1990:APH

[JP90] D. D. Joseph and Luigi Preziosi. Addendum to the paper “Heat waves” [Rev. Mod. Phys. 61, 41 (1989)]. Reviews of Modern Physics, 62(2):375–391, April 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.62.375;   http:/
/rmp.aps.org/abstract/RMP/v62/i2/p375_1. See [JP89].

Jain:1998:NSO

[JSH^{+}98] A. K. Jain, R. K. Sheline, D. M. Headly, P. C. Sood, D. G. Burke, I. Hrivnácová, J. Kvasil, D. Nosek, and R. W. Hoff. Nuclear structure in odd-odd nuclei, 144 < A < 194. Reviews of Modern Physics, 70(3):843–895, July 1998. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.70.843;
 http://rmp.aps.org/abstract/RMP/v70/i3/p843_1.

Jain:1990:ISD

[JSSJ90] A. K. Jain, R. K. Sheline, P. C. Sood, and Kiran Jain. Intrinsic states of deformed odd-A nuclei in the mass regions (151 ≤ A ≤ 193) and (A ≥ 221). Reviews of Modern Physics, 62(2):393–509, April 1990. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.62.393;
 http://rmp.aps.org/abstract/RMP/v62/i2/p393_1.

Julesz:1991:EVF

[Jul91] Bela Julesz. Early vision and focal attention. Reviews of Modern Physics, 63(3):735–772, July 1991. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.63.735;
 http://rmp.aps.org/abstract/RMP/v63/i3/p735_1.

Kadanoff:1999:BUS

[Kad99] Leo P. Kadanoff. Built upon sand: Theoretical ideas inspired by granular flows. Reviews of Modern Physics, 71(1):435–444, January 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.71.435;   http://rmp.aps.org/abstract/
RMP/v71/i1/p435_1.

Kastner:1992:SET

[Kas92] M. A. Kastner. The single-electron transistor. Reviews of Modern Physics, 64(3):849–858, July 1992. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.64.849;
 http://rmp.aps.org/abstract/RMP/v64/i3/p849_1.

Kastner:1998:MTO

[KBSE98] M. A. Kastner, R. J. Birgeneau, G. Shirane, and Y. Endoh. Magnetic, transport, and optical properties of monolayer copper oxides. Reviews of Modern Physics, 70(3):897–928, July 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.70.897;   http://rmp.aps.org/abstract/
RMP/v70/i3/p897_1.

Kendall:1991:DIS

[Ken91] Henry W. Kendall. Deep inelastic scattering: Experiments on the proton and the observation of scaling. Reviews of Modern Physics, 63(3):597–614, July 1991. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.63.597;
 http://rmp.aps.org/abstract/RMP/v63/i3/p597_1.

Kardar:1999:FVO

[KG99] Mehran Kardar and Ramin Golestanian. The “friction” of vacuum, and other fluctuation-induced forces. Reviews of Modern Physics, 71(4):1233–1245, July 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.1233;   http:
//rmp.aps.org/abstract/RMP/v71/i4/p1233_1.

Khitrova:1999:NON

[KGJ^{+}99] G. Khitrova, H. M. Gibbs, F. Jahnke, M. Kira, and S. W. Koch. Nonlinear optics of normal-mode-coupling semiconductor microcavities. Reviews of Modern Physics, 71(5):1591– 1639, October 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.1591;   http://
rmp.aps.org/abstract/RMP/v71/i5/p1591_1.

Kirsten:1999:SNE

[Kir99] Till A. Kirsten. Solar neutrino experiments: results and implications. Reviews of Modern Physics, 71(4):1213–1232, July 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.1213;   http://rmp.aps.org/abstract/RMP/
v71/i4/p1213_1.

Koelle:1999:EHT

[KKL^{+}99a] D. Koelle, R. Kleiner, F. Ludwig, E. Dantsker, and John Clarke. Erratum: High-transition-temperature superconducting quantum interference devices [Rev. Mod. Phys. 71, 631 (1999)]. Reviews of Modern Physics, 71(4):1249, July 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.71.1249;   http:/
/rmp.aps.org/abstract/RMP/v71/i4/p1249_1. See [KKL^{+}99b].

Koelle:1999:HTT

[KKL^{+}99b] D. Koelle, R. Kleiner, F. Ludwig, E. Dantsker, and John Clarke. High-transition-temperature superconducting quantum interference devices. Reviews of Modern Physics, 71(3):631–686, April 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.71.631;   http://rmp.aps.org/abstract/

RMP/v71/i3/p631_1. See erratum [KKL^{+}99a].

Klevansky:1992:NJL

[Kle92] S. P. Klevansky. The Nambu–Jona–Lasinio model of quantum chromodynamics. Reviews of Modern Physics, 64(3):649–708, July 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.64.649;   http://rmp.aps.org/abstract/
RMP/v64/i3/p649_1.

Klein:1999:SBP

[Kle99a] Spencer Klein. Suppression of Bremsstrahlung and pair production due to environmental factors. Reviews of Modern Physics, 71(5):1501–1538, October 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.1501;   http://
rmp.aps.org/abstract/RMP/v71/i5/p1501_1.

Kleppner:1999:SHA

[Kle99b] Daniel Kleppner. A short history of atomic physics in the Twentieth Century. Reviews of Modern Physics, 71(2):S78–S84, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.71.S78;   http://rmp.aps.org/abstract/
RMP/v71/i2/pS78_1.

Kivshar:1989:DSN

[KM89] Yuri S. Kivshar and Boris A. Malomed. Dynamics of solitons in nearly integrable systems. Reviews of Modern Physics, 61 (4):763–915, October 1989. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.61.763;   http:/
/rmp.aps.org/abstract/RMP/v61/i4/p763_1. See addendum

[KM91a].

Kivshar:1991:ADS

[KM91a] Yuri S. Kivshar and Boris A. Malomed. Addendum: Dynamics of solitons in nearly integrable systems. Reviews of Modern Physics, 63(1):211, January 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.63.211;   http:/
/rmp.aps.org/abstract/RMP/v63/i1/p211_1. See [KM89].

Klein:1991:BRL

[KM91b] Abraham Klein and E. R. Marshalek. Boson realizations of Lie algebras with applications to nuclear physics. Reviews of Modern Physics, 63(2):375–558, April 1991. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.63.375;
 http://rmp.aps.org/abstract/RMP/v63/i2/p375_1.

Koch:1994:BPF

[KM94] A. J. Koch and H. Meinhardt. Biological pattern formation: from basic mechanisms to complex structures. Reviews of Modern Physics, 66(4):1481–1507, October 1994. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.66.1481;   http://rmp.aps.org/abstract/RMP/
v66/i4/p1481_1.

Kaganer:1999:SPT

[KMD99] Vladimir M. Kaganer, Helmuth Möhwald, and Pulak Dutta. Structure and phase transitions in Langmuir monolayers. Reviews of Modern Physics, 71(3):779–819, April 1999. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.71.779;
 http://rmp.aps.org/abstract/RMP/v71/i3/p779_1.

Mao:1994:UPT

[kMH94] Ho kwang Mao and Russell J. Hemley. Ultrahigh-pressure transitions in solid hydrogen. Reviews of Modern Physics, 66 (2):671–692, April 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.66.671;   http://
rmp.aps.org/abstract/RMP/v66/i2/p671_1.

Kohn:1999:ECM

[Koh99a] W. Kohn. An essay on condensed matter physics in the twentieth century. Reviews of Modern Physics, 71(2):S59–S77, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.71.S59;   http://rmp.aps.org/abstract/
RMP/v71/i2/pS59_1.

Kohn:1999:NLE

[Koh99b] W. Kohn. Nobel Lecture: Electronic structure of matter-wave functions and density functionals. Reviews of Modern Physics, 71 (5):1253–1266, October 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.1253;   http://
rmp.aps.org/abstract/RMP/v71/i5/p1253_1.

Kolobov:1999:SBN

[Kol99] Mikhail I. Kolobov. The spatial behavior of nonclassical light. Reviews of Modern Physics, 71(5):1539–1589, October 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.1539;   http://rmp.aps.org/abstract/RMP/
v71/i5/p1539_1.

Kroto:1997:SSS

[Kro97] Harold Kroto. Symmetry, space, stars and C_{60}. Reviews of Modern Physics, 69(3):703–722, July 1997. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.69.703;
 http://rmp.aps.org/abstract/RMP/v69/i3/p703_1.

Kolomeisky:1996:PDC

[KS96] Eugene B. Kolomeisky and Joseph P. Straley. Phase diagram and correlation exponents for interacting fermions in one dimension. Reviews of Modern Physics, 68(1):175–214, January 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.68.175;   http://rmp.aps.org/abstract/
RMP/v68/i1/p175_1.

Laughlin:1999:NLF

[Lau99] R. B. Laughlin. Nobel Lecture: Fractional quantization. Reviews of Modern Physics, 71(4):863–874, July 1999. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.71.863;
 http://rmp.aps.org/abstract/RMP/v71/i4/p863_1.

Lynden-Bell:1994:TRC

[LBM94] R. M. Lynden-Bell and K. H. Michel. Translation-rotation coupling, phase transitions, and elastic phenomena in orientationally disordered crystals. Reviews of Modern Physics, 66(3):721–762, July 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.66.721;   http://rmp.aps.org/abstract/
RMP/v66/i3/p721_1.

Lynden-Bell:1998:CMN

[LBNZ98] D. Lynden-Bell and M. Nouri-Zonoz. Classical monopoles: Newton, NUT space, gravomagnetic lensing, and atomic spectra. Reviews of Modern Physics, 70(2):427–445, April 1998. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.70.427;
 http://rmp.aps.org/abstract/RMP/v70/i2/p427_1.

Leggett:1987:DDT

[LCD^{+}87] A. J. Leggett, S. Chakravarty, A. T. Dorsey, Matthew P. A. Fisher, Anupam Garg, and W. Zwerger. Dynamics of the dissipative two-state system. Reviews of Modern Physics, 59(1): 1–85, January 1987. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://link.

aps.org/doi/10.1103/RevModPhys.59.1;   http://rmp.aps.
org/abstract/RMP/v59/i1/p1_1. See erratum [LCD^{+}95].

Leggett:1995:EDD

[LCD^{+}95] Anthony J. Leggett, S. Chakravarty, A. T. Dorsey, Matthew P. A. Fisher, Anupam Garg, and W. Zwerger. Erratum: Dynamics of the dissipative two-state system. Reviews of Modern Physics, 67(3):725–726, July 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.67.725;   http:/
/rmp.aps.org/abstract/RMP/v67/i3/p725_1. See [LCD^{+}87].

Lebowitz:1999:SMS

[Leb99] Joel L. Lebowitz. Statistical mechanics: a selective review of two central issues. Reviews of Modern Physics, 71(2): S346–S357, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S346;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS346_1.

Lee:1997:EPL

[Lee97] David M. Lee. The extraordinary phases of liquid ^{3}He. Reviews of Modern Physics, 69(3):645–666, July 1997. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.69.645;
 http://rmp.aps.org/abstract/RMP/v69/i3/p645_1. See erra-

tum [Lee98].

Lee:1998:EEP

[Lee98] David M. Lee. Erratum: The extraordinary phases of liquid 3He [Rev. Mod. Phys. 69, 645 (1997)]. Reviews of Modern Physics, 70(1):319, January 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.70.319;   http:/
/rmp.aps.org/abstract/RMP/v70/i1/p319_1. See [Lee97].

Leggett:1999:S

[Leg99] A. J. Leggett. Superfluidity. Reviews of Modern Physics, 71 (2):S318–S323, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S318;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS318_1.

Lifshitz:1997:TCS

[Lif97] Ron Lifshitz. Theory of color symmetry for periodic and quasiperiodic crystals. Reviews of Modern Physics, 69(4): 1181–1218, October 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.69.1181;   http://
rmp.aps.org/abstract/RMP/v69/i4/p1181_1.

Lissauer:1999:CMS

[Lis99] Jack J. Lissauer. Chaotic motion in the Solar System. Reviews of Modern Physics, 71(3):835–845, April 1999. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.71.835;
 http://rmp.aps.org/abstract/RMP/v71/i3/p835_1.

Lidsey:1997:RIP

[LLK^{+}97] James E. Lidsey, Andrew R. Liddle, Edward W. Kolb, Edmund J. Copeland, Tiago Barreiro, and Mark Abney. Reconstructing the inflaton potential — an overview. Reviews of Modern Physics, 69(2):373–410, April 1997. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.69.373;
 http://rmp.aps.org/abstract/RMP/v69/i2/p373_1.

Lechanoine-LeLuc:1993:NNE

[LLL93] Catherine Lechanoine-LeLuc and Francois Lehar. Nucleonnucleon elastic scattering and total cross sections. Reviews of Modern Physics, 65(1):47–86, January 1993. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
65.47;   http://rmp.aps.org/abstract/RMP/v65/i1/p47_1.

Langacker:1992:HPE

[LLM92] Paul Langacker, Mingxing Luo, and Alfred K. Mann. Highprecision electroweak experiments: a global search for new physics beyond the Standard Model. Reviews of Modern Physics, 64 (1):87–192, January 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.64.87;   http://
rmp.aps.org/abstract/RMP/v64/i1/p87_1.

Landauer:1994:BIT

[LM94] R. Landauer and Th. Martin. Barrier interaction time in tunneling. Reviews of Modern Physics, 66(1):217–228, January 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.66.217;   http://rmp.aps.org/abstract/
RMP/v66/i1/p217_1.

Littlejohn:1997:GFS

[LR97] Robert G. Littlejohn and Matthias Reinsch. Gauge fields in the separation of rotations andinternal motions in the n-body problem. Reviews of Modern Physics, 69(1):213–276, January 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.69.213;   http://rmp.aps.org/abstract/
RMP/v69/i1/p213_1.

Lamb:1999:LPQ

[LSST99] W. E. Lamb, W. P. Schleich, M. O. Scully, and C. H. Townes. Laser physics: Quantum controversy in action. Reviews of Modern Physics, 71(2):S263–S273, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S263;   http:
//rmp.aps.org/abstract/RMP/v71/i2/pS263_1.

Miller:1991:SWV

[MA91] Gregory H. Miller and Thomas J. Ahrens. Shock-wave viscosity measurement. Reviews of Modern Physics, 63(4):919–948, October 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.63.919;   http://rmp.aps.org/abstract/
RMP/v63/i4/p919_1.

Manousakis:1991:SHA

[Man91] Efstratios Manousakis. The spin-1/2 Heisenberg antiferromagnet on a square lattice and its application to the cuprous oxides. Reviews of Modern Physics, 63(1):1–62, January 1991. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.63.1;
http://rmp.aps.org/abstract/RMP/v63/i1/p1_1.

Mandel:1999:QEO

[Man99] L. Mandel. Quantum effects in one-photon and two-photon interference. Reviews of Modern Physics, 71(2):S274–S282, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.71.S274;   http://rmp.aps.org/abstract/
RMP/v71/i2/pS274_1.

Marcus:1993:ETR

[Mar93] Rudolph A. Marcus. Electron transfer reactions in chemistry. Theory and experiment. Reviews of Modern Physics, 65(3):599–610, July 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.65.599;   http://rmp.aps.org/abstract/
RMP/v65/i3/p599_1.

Myers:1992:HID

[MBB^{+}92] S. M. Myers, M. I. Baskes, H. K. Birnbaum, J. W. Corbett, G. G. DeLeo, S. K. Estreicher, E. E. Haller, P. Jena, N. M. Johnson, R. Kirchheim, S. J. Pearton, and M. J. Stavola. Hydrogen interactions with defects in crystalline solids. Reviews of Modern Physics, 64(2):559–617, April 1992. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.64.559;
 http://rmp.aps.org/abstract/RMP/v64/i2/p559_1.

Mitchell:1999:PVC

[MBW99] G. E. Mitchell, J. D. Bowman, and H. A. Weidenmüller. Parity violation in the compound nucleus. Reviews of Modern Physics, 71(1):445–457, January 1999. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.71.445;
 http://rmp.aps.org/abstract/RMP/v71/i1/p445_1.

Mead:1992:GPM

[Mea92] C. Alden Mead. The geometric phase in molecular systems. Reviews of Modern Physics, 64(1):51–85, January 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
64.51;   http://rmp.aps.org/abstract/RMP/v64/i1/p51_1.

Meerson:1996:NDR

[Mee96] Baruch Meerson. Nonlinear dynamics of radiative condensations in optically thin plasmas. Reviews of Modern Physics, 68 (1):215–257, January 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.68.215;   http://
rmp.aps.org/abstract/RMP/v68/i1/p215_1.

Meiss:1992:SMV

[Mei92] J. D. Meiss. Symplectic maps, variational principles, and transport. Reviews of Modern Physics, 64(3):795–848, July 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.64.795;   http://rmp.aps.org/abstract/
RMP/v64/i3/p795_1.

Mermin:1992:ESGa

[Mer92a] N. David Mermin. Erratum: The space groups of icosahedral quasicrystals and cubic, orthorhombic, monoclinic, and triclinic crystals. Reviews of Modern Physics, 64(2):635, April 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
64.635;   http://rmp.aps.org/abstract/RMP/v64/i2/p635_

1. See [Mer92c, Mer92b, Mer94].

Mermin:1992:ESGb

[Mer92b] N. David Mermin. Erratum: The space groups of icosahedral quasicrystals and cubic, orthorhombic, monoclinic, and triclinic crystals. Reviews of Modern Physics, 64(4):1163, October 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.64.1163;   http://rmp.aps.org/abstract/RMP/

v64/i4/p1163_1. See [Mer92c, Mer92a, Mer94].

Mermin:1992:SGI

[Mer92c] N. David Mermin. The space groups of icosahedral quasicrystals and cubic, orthorhombic, monoclinic, and triclinic crystals. Reviews of Modern Physics, 64(1):3–49, January 1992. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.64.3;
http://rmp.aps.org/abstract/RMP/v64/i1/p3_1. See errata

[Mer92a, Mer92b, Mer94].

Mermin:1993:HVT

[Mer93] N. David Mermin. Hidden variables and the two theorems of John Bell. Reviews of Modern Physics, 65(3):803–815, July 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.65.803;   http://rmp.aps.org/abstract/

RMP/v65/i3/p803_1. See erratum [Mer16].

Mermin:1994:ESG

[Mer94] N. David Mermin. Erratum: The space groups of icosahedral quasicrystals and cubic, orthorhombic, monoclinic, and triclinic crystals. Reviews of Modern Physics, 66(1):249, January 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.66.249;   http://rmp.aps.org/abstract/

RMP/v66/i1/p249_1. See [Mer92c, Mer92a, Mer92b].

Mermin:2016:EHV

[Mer16] N. David Mermin. Erratum: Hidden variables and the two theorems of John Bell [Rev. Mod. Phys. 65, 803 (1993)]. Reviews of Modern Physics, 88(??):039902–??, March 2016. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://journals.aps.org/rmp/abstract/10.1103/
RevModPhys.88.039902. See [Mer93].

Mattis:1998:UQF

[MG98] Daniel C. Mattis and M. Lawrence Glasser. The uses of quantum field theory in diffusion-limited reactions. Reviews of Modern Physics, 70(3):979–1001, July 1998. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.70.979;
 http://rmp.aps.org/abstract/RMP/v70/i3/p979_1.

Martinez:1999:PTE

[MMRT99] M. Martinez, R. Miquel, L. Rolandi, and R. Tenchini. Precision tests of the electroweak interaction at the Z pole. Reviews of Modern Physics, 71(3):575–629, April 1999. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.71.575;
 http://rmp.aps.org/abstract/RMP/v71/i3/p575_1.

Marsili:1996:SGE

[MMTB96] Matteo Marsili, Amos Maritan, Flavio Toigo, and Jayanth R. Banavar. Stochastic growth equations and reparametrization invariance. Reviews of Modern Physics, 68(4):963–983, October 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.68.963;   http://rmp.aps.org/abstract/
RMP/v68/i4/p963_1.

Meyer-Ortmanns:1996:PTQ

[MO96] Hildegard Meyer-Ortmanns. Phase transitions in quantum chromodynamics. Reviews of Modern Physics, 68(2):473–598, April 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.68.473;   http://rmp.aps.org/abstract/
RMP/v68/i2/p473_1.

Moon:1993:ACT

[Moo93] Hie-Tae Moon. Approach to chaos through instabilities. Reviews of Modern Physics, 65(4):1535–1543, October 1993. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.65.1535;   http://rmp.aps.org/abstract/RMP/
v65/i4/p1535_1.

Morrison:1998:HDI

[Mor98] P. J. Morrison. Hamiltonian description of the ideal fluid. Reviews of Modern Physics, 70(2):467–521, April 1998. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.70.467;
 http://rmp.aps.org/abstract/RMP/v70/i2/p467_1.

Merzhanov:1999:PRW

[MR99] A. G. Merzhanov and E. N. Rumanov. Physics of reaction waves. Reviews of Modern Physics, 71(4):1173–1211, July 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.1173;   http://rmp.aps.org/abstract/RMP/
v71/i4/p1173_1.

Micnas:1990:SNB

[MRR90] R. Micnas, J. Ranninger, and S. Robaszkiewicz. Superconductivity in narrow-band systems with local nonretarded attractive interactions. Reviews of Modern Physics, 62(1):113–171, January 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.62.113;   http://rmp.aps.org/abstract/
RMP/v62/i1/p113_1.

Nagel:1992:IS

[Nag92] Sidney R. Nagel. Instabilities in a sandpile. Reviews of Modern Physics, 64(1):321–325, January 1992. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.64.321;
 http://rmp.aps.org/abstract/RMP/v64/i1/p321_1.

Nenciu:1991:DBE

[Nen91] G. Nenciu. Dynamics of band electrons in electric and magnetic fields: rigorous justification of the effective Hamiltonians. Reviews of Modern Physics, 63(1):91–127, January 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
63.91;   http://rmp.aps.org/abstract/RMP/v63/i1/p91_1.

Nemirovskii:1995:CQV

[NF95] S. K. Nemirovskii and W. Fiszdon. Chaotic quantized vortices and hydrodynamic processes in superfluid helium. Reviews of Modern Physics, 67(1):37–84, January 1995. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
67.37;   http://rmp.aps.org/abstract/RMP/v67/i1/p37_1.

Nakayama:1994:DPF

[NYO94] Tsuneyoshi Nakayama, Kousuke Yakubo, and Raymond L. Orbach. Dynamical properties of fractal networks: Scaling, numerical simulations, and physical realizations. Reviews of Modern Physics, 66(2):381–443, April 1994. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.66.381;
 http://rmp.aps.org/abstract/RMP/v66/i2/p381_1.

ONeil:1999:CNH

[OC99] T. M. O’Neil and F. V. Coroniti. The collisionless nature of high-temperature plasmas. Reviews of Modern Physics, 71 (2):S404–S410, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S404;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS404_1.

Oron:1997:LSE

[ODB97] Alexander Oron, Stephen H. Davis, and S. George Bankoff. Longscale evolution of thin liquid films. Reviews of Modern Physics, 69(3):931–980, July 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.69.931;   http://
rmp.aps.org/abstract/RMP/v69/i3/p931_1.

Oka:1992:ISH

[Oka92] Takeshi Oka. The infrared spectrum of H_{3}^{+} in laboratory and space plasmas. Reviews of Modern Physics, 64(4):1141– 1149, October 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.64.1141;   http://
rmp.aps.org/abstract/RMP/v64/i4/p1141_1.

Oja:1997:NMO

[OL97] A. S. Oja and O. V. Lounasmaa. Nuclear magnetic ordering in simple metals at positive and negative nanokelvin temperatures. Reviews of Modern Physics, 69(1):1–136, January 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
69.1;   http://rmp.aps.org/abstract/RMP/v69/i1/p1_1.

Omnes:1992:CIQ

[Omn92] Roland Omnès. Consistent interpretations of quantum mechanics. Reviews of Modern Physics, 64(2):339–382, April 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
64.339;   http://rmp.aps.org/abstract/RMP/v64/i2/p339_
1.

Osheroff:1997:SDU

[Osh97] Douglas D. Osheroff. Superfluidity in ^{3}He: Discovery and understanding. Reviews of Modern Physics, 69(3):667–682, July 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.69.667;   http://rmp.aps.org/abstract/
RMP/v69/i3/p667_1.

Osterfeld:1992:NSI

[Ost92] Franz Osterfeld. Nuclear spin and isospin excitations. Reviews of Modern Physics, 64(2):491–557, April 1992. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.64.491;
 http://rmp.aps.org/abstract/RMP/v64/i2/p491_1.

Ostro:1993:PRA

[Ost93]  Steven J. Ostro.  Planetary radar astronomy.  Reviews

of Modern Physics, 65(4):1235–1279, October 1993. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.65.1235;   http://rmp.aps.org/abstract/RMP/
v65/i4/p1235_1.

Ohtaka:1990:TSX

[OT90] K. Ohtaka and Y. Tanabe. Theory of the soft-X-ray edge problem in simple metals: historical survey and recent developments. Reviews of Modern Physics, 62(4):929–991, October 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
62.929;   http://rmp.aps.org/abstract/RMP/v62/i4/p929_
1.

Packard:1998:RJA

[Pac98] Richard E. Packard. The role of the Josephson–Anderson equation in superfluid helium. Reviews of Modern Physics, 70 (2):641–651, April 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.70.641;   http://
rmp.aps.org/abstract/RMP/v70/i2/p641_1.

Pais:1999:TPP

[Pai99] A. Pais. Theoretical particle physics. Reviews of Modern Physics, 71(2):S16–S24, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.71.S16;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS16_1.

Paul:1990:ETC

[Pau90] Wolfgang Paul. Electromagnetic traps for charged and neutral particles. Reviews of Modern Physics, 62(3):531–540, July 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.62.531;   http://rmp.aps.org/abstract/
RMP/v62/i3/p531_1.

Panofsky:1999:AD

[PB99] W. K. H. Panofsky and M. Breidenbach. Accelerators and detectors. Reviews of Modern Physics, 71(2):S121–S132, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S121;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS121_1.

Peierls:1993:EWS

[Pei93] Rudolf Peierls. Early work on solids, mainly in the thirties. Reviews of Modern Physics, 65(2):251–254, April 1993. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.65.251;
 http://rmp.aps.org/abstract/RMP/v65/i2/p251_1.

Pethick:1992:CNS

[Pet92] C. J. Pethick. Cooling of neutron stars. Reviews of Modern Physics, 64(4):1133–1140, October 1992. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
64.1133;   http://rmp.aps.org/abstract/RMP/v64/i4/p1133_
1.

Pfeifer:1995:GTE

[PF95] Peter Pfeifer and Jürg Fröhlich. Generalized time-energy uncertainty relations and bounds on lifetimes of resonances. Reviews of Modern Physics, 67(4):759–779, October 1995. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.67.759;
 http://rmp.aps.org/abstract/RMP/v67/i4/p759_1.

Phillips:1998:NLL

[Phi98] William D. Phillips. Nobel Lecture: Laser cooling and trapping of neutral atoms. Reviews of Modern Physics, 70(3):721–741, July 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.70.721;   http://rmp.aps.org/abstract/
RMP/v70/i3/p721_1.

Pines:1992:ERC

[Pin92] David Pines. Editorial: RMP Colloquia. Reviews of Modern Physics, 64(1):1, January 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.64.1;
http://rmp.aps.org/abstract/RMP/v64/i1/p1_1.

Plenio:1998:QJA

[PK98] M. B. Plenio and P. L. Knight. The quantum-jump approach to dissipative dynamics in quantum optics. Reviews of Modern Physics, 70(1):101–144, January 1998. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.70.101;
 http://rmp.aps.org/abstract/RMP/v70/i1/p101_1.

Puska:1994:TPS

[PN94] M. J. Puska and R. M. Nieminen. Theory of positrons in solids and on solid surfaces. Reviews of Modern Physics, 66(3):841–897, July 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.66.841;   http://rmp.aps.org/abstract/
RMP/v66/i3/p841_1.

Polchinski:1996:SD

[Pol96] Joseph Polchinski. String duality. Reviews of Modern Physics, 68(4):1245–1258, October 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.68.1245;   http://
rmp.aps.org/abstract/RMP/v68/i4/p1245_1.

Pople:1999:NLQ

[Pop99] John A. Pople. Nobel Lecture: Quantum chemical models. Reviews of Modern Physics, 71(5):1267–1274, October 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.1267;   http://rmp.aps.org/abstract/RMP/
v71/i5/p1267_1.

Pound:1999:RNM

[Pou99] Robert V. Pound. From radar to nuclear magnetic resonance. Reviews of Modern Physics, 71(2):S54–S58, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
71.S54;   http://rmp.aps.org/abstract/RMP/v71/i2/pS54_
1.

Palmer:1992:RES

[PR92] R. E. Palmer and P. J. Rous. Resonances in electron scattering by molecules on surfaces. Reviews of Modern Physics, 64(2):383–440, April 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.64.383;   http://
rmp.aps.org/abstract/RMP/v64/i2/p383_1.

Pai:1993:PE

[PS93] Damodar M. Pai and B. E. Springett. Physics of electrophotography. Reviews of Modern Physics, 65(1):163–211, January 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.65.163;   http://rmp.aps.org/abstract/
RMP/v65/i1/p163_1.

Pennington:1996:NMR

[PS96] Charles H. Pennington and Victor A. Stenger. Nuclear magnetic resonance of C_{60} and fulleride superconductors. Reviews of Modern Physics, 68(3):855–910, July 1996. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.68.855;
 http://rmp.aps.org/abstract/RMP/v68/i3/p855_1.

Pandharipande:1997:IPM

[PSdH97] Vijay R. Pandharipande, Ingo Sick, and Peter K. A. deWitt Huberts. Independent particle motion and correlations in fermion systems. Reviews of Modern Physics, 69(3):981–991, July 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.69.981;   http://rmp.aps.org/abstract/
RMP/v69/i3/p981_1.

Payne:1992:IMT

[PTA^{+}92] M. C. Payne, M. P. Teter, D. C. Allan, T. A. Arias, and J. D. Joannopoulos. Iterative minimization techniques for ab initio total-energy calculations: molecular dynamics and conjugate gradients. Reviews of Modern Physics, 64(4):1045– 1097, October 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.64.1045;   http://
rmp.aps.org/abstract/RMP/v64/i4/p1045_1.

Perelson:1997:IP

[PW97] Alan S. Perelson and Gérard Weisbuch. Immunology for physicists. Reviews of Modern Physics, 69(4):1219–1268, October 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.69.1219;   http://rmp.aps.org/abstract/RMP/
v69/i4/p1219_1.

Page:1999:CMB

[PW99] Lyman Page and David Wilkinson. Cosmic microwave background radiation. Reviews of Modern Physics, 71(2):S173– S179, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S173;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS173_1.

Quilichini:1997:PEQ

[Qui97] M. Quilichini. Phonon excitations in quasicrystals. Reviews of Modern Physics, 69(1):277–314, January 1997. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.69.277;
 http://rmp.aps.org/abstract/RMP/v69/i1/p277_1.

Ramsey:1990:ESO

[Ram90] Norman F. Ramsey. Experiments with separated oscillatory fields and hydrogen masers. Reviews of Modern Physics, 62(3):541–552, July 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.62.541;   http://rmp.aps.org/abstract/
RMP/v62/i3/p541_1.

Rammer:1991:QTT

[Ram91] Jørgen Rammer. Quantum transport theory of electrons in solids: a single-particle approach. Reviews of Modern Physics, 63(4):781–817, October 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.63.781;   http://
rmp.aps.org/abstract/RMP/v63/i4/p781_1.

Rau:1992:ARM

[Rau92] A. R. P. Rau. The asymmetric rotor as a model for localization. Reviews of Modern Physics, 64(2):623–632, April 1992. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.64.623;
 http://rmp.aps.org/abstract/RMP/v64/i2/p623_1.

Richman:1995:LSD

[RB95] Jeffrey D. Richman and Patricia R. Burchat. Leptonic and semileptonic decays of charm and bottom hadrons. Reviews of Modern Physics, 67(4):893–976, October 1995. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.67.893;
 http://rmp.aps.org/abstract/RMP/v67/i4/p893_1.

Rieke:1998:SPD

[RB98] F. Rieke and D. A. Baylor. Single-photon detection by rod cells of the retina. Reviews of Modern Physics, 70(3): 1027–1036, July 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.70.1027;   http://
rmp.aps.org/abstract/RMP/v70/i3/p1027_1.

Reeves:1994:OLE

[Ree94] Hubert Reeves. On the origin of the light elements (Z < 6). Reviews of Modern Physics, 66(1):193–216, January 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
66.193;   http://rmp.aps.org/abstract/RMP/v66/i1/p193_
1.

Reines:1996:NPP

[Rei96] F. Reines. The neutrino: from poltergeist to particle. Reviews of Modern Physics, 68(2):317–327, April 1996. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.68.317;
 http://rmp.aps.org/abstract/RMP/v68/i2/p317_1.

Resta:1994:MPC

[Res94] Raffaele Resta. Macroscopic polarization in crystalline dielectrics: the geometric phase approach. Reviews of Modern Physics, 66(3):899–915, July 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.66.899;   http://
rmp.aps.org/abstract/RMP/v66/i3/p899_1.

Riordan:1999:IT

[RHH99] Michael Riordan, Lillian Hoddeson, and Conyers Herring. The invention of the transistor. Reviews of Modern Physics, 71 (2):S336–S345, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S336;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS336_1.

Richardson:1997:PE

[Ric97] Robert C. Richardson. The Pomeranchuk effect. Reviews of Modern Physics, 69(3):683–690, July 1997. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.69.683;
 http://rmp.aps.org/abstract/RMP/v69/i3/p683_1.

Riisager:1994:NHS

[Rii94] K. Riisager. Nuclear halo states. Reviews of Modern Physics, 66(3):1105–1116, July 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.66.1105;   http://
rmp.aps.org/abstract/RMP/v66/i3/p1105_1.

Rudd:1992:EPP

[RKMG92] M. E. Rudd, Y.-K. Kim, D. H. Madison, and T. J. Gay. Electron production in proton collisions with atoms and molecules: energy distributions. Reviews of Modern Physics, 64(2):441–490, April 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.64.441;   http://rmp.aps.org/abstract/
RMP/v64/i2/p441_1.

Rabson:1991:SGA

[RMRW91] David A. Rabson, N. David Mermin, Daniel S. Rokhsar, and David C. Wright. The space groups of axial crystals and quasicrystals. Reviews of Modern Physics, 63(3):699–733, July 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.63.699;   http://rmp.aps.org/abstract/
RMP/v63/i3/p699_1.

Robertson:1991:GAV

[Rob91] Douglas S. Robertson. Geophysical applications of very-long-baseline interferometry. Reviews of Modern Physics, 63(4): 899–918, October 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.63.899;   http://
rmp.aps.org/abstract/RMP/v63/i4/p899_1.

Roberts:1992:BHE

[Rob92] Arthur Roberts. The birth of high-energy neutrino astronomy: a personal history of the DUMAND project. Reviews of Modern Physics, 64(1):259–312, January 1992. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.64.259;
 http://rmp.aps.org/abstract/RMP/v64/i1/p259_1.

Robinson:1997:NWC

[Rob97] P. A. Robinson. Nonlinear wave collapse and strong turbulence. Reviews of Modern Physics, 69(2):507–574, April 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
69.507;   http://rmp.aps.org/abstract/RMP/v69/i2/p507_
1.

Rosner:1992:EMT

[Ros92] Jonathan L. Rosner. Electroweak measurements and the top quark. Reviews of Modern Physics, 64(4):1151–1161, October 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.64.1151;   http://rmp.aps.org/abstract/

RMP/v64/i4/p1151_1. See erratum [Ros94].

Rosner:1994:EEM

[Ros94] Jonathan L. Rosner. Erratum: Electroweak measurements and the top quark. Reviews of Modern Physics, 66(3):1117, July 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.66.1117;   http://rmp.aps.org/abstract/RMP/
v66/i3/p1117_1. See [Ros92].

Rasolt:1992:TAS

[RT92] Mark Rasolt and Zlatko Tešanović. Theoretical aspects of superconductivity in very high magnetic fields. Reviews of Modern Physics, 64(3):709–754, July 1992. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.64.709;
 http://rmp.aps.org/abstract/RMP/v64/i3/p709_1.

Ruppeiner:1995:RGT

[Rup95] George Ruppeiner. Riemannian geometry in thermodynamic fluctuation theory. Reviews of Modern Physics, 67(3):605–659, July 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.67.605;   http://rmp.aps.org/abstract/

RMP/v67/i3/p605_1. See erratum [Rup96].

Ruppeiner:1996:ERG

[Rup96] George Ruppeiner. Erratum: Riemannian geometry in thermodynamic fluctuation theory. Reviews of Modern Physics, 68(1):313, January 1996. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.68.313;   http://
rmp.aps.org/abstract/RMP/v68/i1/p313_1. See [Rup95].

Ritchie:1993:RKD

[RW93] Jack L. Ritchie and Stanley G. Wojcicki. Rare K decays. Reviews of Modern Physics, 65(4):1149–1197, October 1993. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.65.1149;   http://rmp.aps.org/abstract/RMP/
v65/i4/p1149_1.

Roggemann:1997:IRG

[RWF97] Michael C. Roggemann, Byron M. Welsh, and Robert Q. Fugate. Improving the resolution of ground-based telescopes. Reviews of Modern Physics, 69(2):437–506, April 1997. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.69.437;
 http://rmp.aps.org/abstract/RMP/v69/i2/p437_1.

Rothman:1994:LGM

[RZ94] Daniel H. Rothman and Stéphane Zaleski. Lattice-gas models of phase separation: interfaces, phase transitions, and multiphase flow. Reviews of Modern Physics, 66(4):1417–1479, October 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.66.1417;   http://rmp.aps.org/abstract/RMP/
v66/i4/p1417_1.

Sadoulet:1999:DND

[Sad99] Bernard Sadoulet. Deciphering the nature of dark matter. Reviews of Modern Physics, 71(2):S197–S204, March 1999. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S197;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS197_1.

Sahimi:1993:FPR

[Sah93] Muhammad Sahimi. Flow phenomena in rocks: from continuum models to fractals, percolation, cellular automata, and simulated annealing. Reviews of Modern Physics, 65(4):1393– 1534, October 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.65.1393;   http://
rmp.aps.org/abstract/RMP/v65/i4/p1393_1.

Salpeter:1999:SN

[Sal99] Edwin E. Salpeter. Stellar nucleosynthesis. Reviews of Modern Physics, 71(2):S220–S222, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S220;   http:
//rmp.aps.org/abstract/RMP/v71/i2/pS220_1.

Sapirstein:1998:TMR

[Sap98] J. Sapirstein. Theoretical methods for the relativistic atomic many-body problem. Reviews of Modern Physics, 70(1): 55–76, January 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.70.55;   http://
rmp.aps.org/abstract/RMP/v70/i1/p55_1.

Shchukin:1999:SON

[SB99] Vitaliy A. Shchukin and Dieter Bimberg. Spontaneous ordering of nanostructures on crystal surfaces. Reviews of Modern Physics, 71(4):1125–1171, July 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.1125;   http://
rmp.aps.org/abstract/RMP/v71/i4/p1125_1.

Segel:1991:IIM

[Seg91] L. A. Segel. The infinite and the infinitesimal in models for natural phenomena. Reviews of Modern Physics, 63(2):225–238, April 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.63.225;   http://rmp.aps.org/abstract/
RMP/v63/i2/p225_1.

Sovers:1998:AGR

[SFJ98] Ojars J. Sovers, John L. Fanselow, and Christopher S. Jacobs. Astrometry and geodesy with radio interferometry: experiments, models, results. Reviews of Modern Physics, 70(4): 1393–1454, October 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.70.1393;   http://
rmp.aps.org/abstract/RMP/v70/i4/p1393_1.

Sidles:1995:MRF

[SGB^{+}95] J. A. Sidles, J. L. Garbini, K. J. Bruland, D. Rugar, O. Züger, S. Hoen, and C. S. Yannoni. Magnetic resonance force microscopy. Reviews of Modern Physics, 67(1):249–265, January 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.67.249;   http://rmp.aps.org/abstract/
RMP/v67/i1/p249_1.

Sondhi:1997:CQP

[SGCS97] S. L. Sondhi, S. M. Girvin, J. P. Carini, and D. Shahar. Continuous quantum phase transitions. Reviews of Modern Physics, 69(1):315–333, January 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.69.315;   http://
rmp.aps.org/abstract/RMP/v69/i1/p315_1.

Shankar:1994:RGA

[Sha94] R. Shankar. Renormalization-group approach to interacting fermions. Reviews of Modern Physics, 66(1):129–192, January 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.66.129;   http://rmp.aps.org/abstract/
RMP/v66/i1/p129_1.

Shapiro:1999:CR

[Sha99] Irwin I. Shapiro. A century of Relativity. Reviews of Modern Physics, 71(2):S41–S53, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S41;   http:/
/rmp.aps.org/abstract/RMP/v71/i2/pS41_1.

Sheffield:1994:PMF

[She94] John Sheffield. The physics of magnetic fusion reactors. Reviews of Modern Physics, 66(3):1015–1103, July 1994. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.66.1015;   http://rmp.aps.org/abstract/RMP/

v66/i3/p1015_1. See erratum [She95].

Sheffield:1995:EPM

[She95] John Sheffield. Erratum: The physics of magnetic fusion reactors. Reviews of Modern Physics, 67(1):267, January 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
67.267;   http://rmp.aps.org/abstract/RMP/v67/i1/p267_

1. See [She94].

Shuryak:1993:CFQ

[Shu93] Edward V. Shuryak. Correlation functions in the QCD vacuum. Reviews of Modern Physics, 65(1):1–46, January 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
65.1;   http://rmp.aps.org/abstract/RMP/v65/i1/p1_1.

Shull:1995:EDN

[Shu95] Clifford G. Shull. Early development of neutron scattering. Reviews of Modern Physics, 67(4):753–757, October 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
67.753;   http://rmp.aps.org/abstract/RMP/v67/i4/p753_

Siska:1993:MBS

[Sis93] P. E. Siska. Molecular-beam studies of Penning ionization. Reviews of Modern Physics, 65(2):337–412, April 1993. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.65.337;
 http://rmp.aps.org/abstract/RMP/v65/i2/p337_1.

Slusher:1999:LT

[Slu99] R. E. Slusher. Laser technology. Reviews of Modern Physics, 71(2):S471–S479, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S471;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS471_1.

Smith:1990:TSS

[SM90] D. L. Smith and C. Mailhiot. Theory of semiconductor superlattice electronic structure. Reviews of Modern Physics, 62 (1):173–234, January 1990. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.62.173;   http://
rmp.aps.org/abstract/RMP/v62/i1/p173_1.

Senatore:1994:RPF

[SM94] G. Senatore and N. H. March. Recent progress in the field of electron correlation. Reviews of Modern Physics, 66(2):445–479, April 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.66.445;   http://rmp.aps.org/abstract/
RMP/v66/i2/p445_1.

Smalley:1997:DF

[Sma97] Richard E. Smalley. Discovering the fullerenes. Reviews of Modern Physics, 69(3):723–730, July 1997. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.69.723;
 http://rmp.aps.org/abstract/RMP/v69/i3/p723_1.

Schiffer:1995:NTS

[SO95] P. Schiffer and D. D. Osheroff. Nucleation of the AB transition in superfluid ^{3}He: Surface effects and baked Alaska. Reviews of Modern Physics, 67(2):491–501, April 1995. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.67.491;
 http://rmp.aps.org/abstract/RMP/v67/i2/p491_1.

Storchak:1998:QDM

[SP98] Vyacheslav G. Storchak and Nikolai V. Prokof’ev. Quantum diffusion of muons and muonium atoms in solids. Reviews of Modern Physics, 70(3):929–978, July 1998. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.70.929;
 http://rmp.aps.org/abstract/RMP/v70/i3/p929_1.

Spruch:1991:PNT

[Spr91] Larry Spruch. Pedagogic notes on Thomas–Fermi theory (and on some improvements): atoms, stars, and the stability of bulk matter. Reviews of Modern Physics, 63(1):151–209, January 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.63.151;   http://rmp.aps.org/abstract/
RMP/v63/i1/p151_1.

Sigrist:1995:UPP

[SR95] Manfred Sigrist and T. M. Rice. Unusual paramagnetic phenomena in granular high-temperature superconductors — a consequence of d-wave pairing? Reviews of Modern Physics, 67 (2):503–513, April 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.67.503;   http://
rmp.aps.org/abstract/RMP/v67/i2/p503_1.

Sreenivasan:1999:FT

[Sre99] Katepalli R. Sreenivasan. Fluid turbulence. Reviews of Modern Physics, 71(2):S383–S395, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S383;   http:
//rmp.aps.org/abstract/RMP/v71/i2/pS383_1.

Stolovitzky:1994:KRS

[SS94] G. Stolovitzky and K. R. Sreenivasan. Kolmogorov’s refined similarity hypotheses for turbulence and general stochastic processes. Reviews of Modern Physics, 66(1):229–240, January 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.

1103/RevModPhys.66.229;   http://rmp.aps.org/abstract/
RMP/v66/i1/p229_1.

Surguladze:1996:DWT

[SS96] Levan R. Surguladze and Mark A. Samuel. Decay widths and total cross sections in perturbative QCD. Reviews of Modern Physics, 68(1):259–303, January 1996. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.68.259;
 http://rmp.aps.org/abstract/RMP/v68/i1/p259_1.

Schafer:1998:IQ

[SS98] T. Schäfer and E. V. Shuryak. Instantons in QCD. Reviews of Modern Physics, 70(2):323–425, April 1998. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.70.323;
 http://rmp.aps.org/abstract/RMP/v70/i2/p323_1.

Schwarz:1999:STS

[SS99] John H. Schwarz and Nathan Seiberg. String theory, supersymmetry, unification, and all that. Reviews of Modern Physics, 71 (2):S112–S120, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S112;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS112_1.

Sterman:1995:HPQ

[SSC^{+}95] George Sterman, John Smith, John C. Collins, James Whitmore, Raymond Brock, Joey Huston, Jon Pumplin, Wu-Ki Tung, Hendrik Weerts, Chien-Peng Yuan, Stephen Kuhlmann, Sanjib Mishra, Jorge G. Morfín, Fredrick Olness, Joseph Owens, Jianwei Qiu, and Davison E. Soper. Handbook of perturbative QCD. Reviews of Modern Physics, 67(1):157–248, January 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
67.157;   http://rmp.aps.org/abstract/RMP/v67/i1/p157_
1.

Schnatterly:1992:LFS

[ST92] S. E. Schnatterly and C. Tarrio. Local fields in solids: microscopic aspects for dielectrics. Reviews of Modern Physics, 64(2):619–622, April 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.64.619;   http://
rmp.aps.org/abstract/RMP/v64/i2/p619_1.

Schramm:1998:BBN

[ST98] David N. Schramm and Michael S. Turner. Big-bang nucleosynthesis enters the precision era. Reviews of Modern Physics, 70 (1):303–318, January 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.70.303;   http://
rmp.aps.org/abstract/RMP/v70/i1/p303_1.

Schrieffer:1999:S

[ST99] J. R. Schrieffer and M. Tinkham. Superconductivity. Reviews of Modern Physics, 71(2):S313–S317, March 1999. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S313;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS313_1.

Stanley:1999:SUR

[Sta99] H. Eugene Stanley. Scaling, universality, and renormalization: Three pillars of modern critical phenomena. Reviews of Modern Physics, 71(2):S358–S366, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S358;   http:
//rmp.aps.org/abstract/RMP/v71/i2/pS358_1.

Stormer:1999:FQH

[STG99] Horst L. Stormer, Daniel C. Tsui, and Arthur C. Gossard. The fractional quantum Hall effect. Reviews of Modern Physics, 71 (2):S298–S305, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S298;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS298_1.

Stormer:1999:NLF

[Sto99] Horst L. Stormer. Nobel Lecture: The fractional quantum Hall effect. Reviews of Modern Physics, 71(4):875–889, July 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.71.875;   http://rmp.aps.org/abstract/
RMP/v71/i4/p875_1.

Sigrist:1991:PTU

[SU91] Manfred Sigrist and Kazuo Ueda. Phenomenological theory of unconventional superconductivity. Reviews of Modern Physics, 63(2):239–311, April 1991. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.63.239;   http://
rmp.aps.org/abstract/RMP/v63/i2/p239_1.

Sutherland:1999:GMR

[Sut99] Will Sutherland. Gravitational microlensing — a report on the MACHO project. Reviews of Modern Physics, 71(1):421–434, January 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.71.421;   http://rmp.aps.org/abstract/
RMP/v71/i1/p421_1.

Swanson:1995:CAE

[Swa95] D. G. Swanson. Cyclotron absorption and emission in mode conversion layers — a new paradigm. Reviews of Modern Physics, 67(4):837–862, October 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.67.837;   http://
rmp.aps.org/abstract/RMP/v67/i4/p837_1.

Taylor:1991:DIS

[Tay91] Richard E. Taylor. Deep inelastic scattering: The early years. Reviews of Modern Physics, 63(3):573–595, July 1991. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.63.573;
 http://rmp.aps.org/abstract/RMP/v63/i3/p573_1.

Taylor:1994:BPR

[Tay94] Joseph H. Taylor, Jr. Binary pulsars and relativistic gravity. Reviews of Modern Physics, 66(3):711–719, July 1994. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.66.711;
 http://rmp.aps.org/abstract/RMP/v66/i3/p711_1.

Tsoi:1999:SCE

[TBW99] V. S. Tsoi, J. Bass, and P. Wyder. Studying conductionelectron/interface interactions using transverse electron focusing. Reviews of Modern Physics, 71(5):1641–1693, October 1999.

CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.1641;   http://rmp.aps.org/abstract/RMP/
v71/i5/p1641_1.

Till:1999:NFR

[Til99] Charles E. Till. Nuclear fission reactors. Reviews of Modern Physics, 71(2):S451–S455, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S451;   http:
//rmp.aps.org/abstract/RMP/v71/i2/pS451_1.

Trodden:1999:EB

[Tro99] Mark Trodden. Electroweak baryogenesis. Reviews of Modern Physics, 71(5):1463–1500, October 1999. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
71.1463;   http://rmp.aps.org/abstract/RMP/v71/i5/p1463_
1.

Tsunetsugu:1997:GSP

[TSU97] Hirokazu Tsunetsugu, Manfred Sigrist, and Kazuo Ueda. The ground-state phase diagram of the one-dimensional Kondo lattice model. Reviews of Modern Physics, 69(3):809–864, July 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.69.809;   http://rmp.aps.org/abstract/
RMP/v69/i3/p809_1.

Tsui:1999:NLI

[Tsu99] Daniel C. Tsui. Nobel Lecture: Interplay of disorder and interaction in two-dimensional electron gas in intense magnetic fields. Reviews of Modern Physics, 71(4):891–895, July 1999. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.71.891;
 http://rmp.aps.org/abstract/RMP/v71/i4/p891_1.

Turner:1999:CM

[TT99] Michael S. Turner and J. Anthony Tyson. Cosmology at the millennium. Reviews of Modern Physics, 71(2):S145–S164, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S145;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS145_1.

VanMieghem:1992:TBT

[Van92] Piet Van Mieghem. Theory of band tails in heavily doped semiconductors. Reviews of Modern Physics, 64(3):755–793, July 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.64.755;   http://rmp.aps.org/abstract/
RMP/v64/i3/p755_1.

VanHarlingen:1995:PST

[Van95] D. J. Van Harlingen. Phase-sensitive tests of the symmetry of the pairing state in the high-temperature superconductors — evidence for d_{x^{2}−y^{2}} symmetry. Reviews of Modern Physics, 67(2):515–535, April 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.67.515;   http://
rmp.aps.org/abstract/RMP/v67/i2/p515_1.

Vavrycuk:1994:ENF

[Vav94] Václav Vavrycuk. Elastic near-field wave energy radiated by a spherical cavity. Reviews of Modern Physics, 66(1):241–247, January 1994. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.66.241;   http://rmp.aps.org/abstract/
RMP/v66/i1/p241_1.

Vugmeister:1990:DGF

[VG90] B. E. Vugmeister and M. D. Glinchuk. Dipole glass and ferroelectricity in random-site electric dipole systems. Reviews of Modern Physics, 62(4):993–1026, October 1990. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.62.993;
 http://rmp.aps.org/abstract/RMP/v62/i4/p993_1.

Vos:1995:OEM

[VM95] M. Vos and I. E. McCarthy. Observing electron motion in solids. Reviews of Modern Physics, 67(3):713–723, July 1995. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
67.713;   http://rmp.aps.org/abstract/RMP/v67/i3/p713_

vanRossum:1999:MSC

[vRN99] M. C. W. van Rossum and Th. M. Nieuwenhuizen. Multiple scattering of classical waves: microscopy, mesoscopy, and diffusion. Reviews of Modern Physics, 71(1):313–371, January 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
71.313;   http://rmp.aps.org/abstract/RMP/v71/i1/p313_
1.

Weiner:1999:ETC

[WBZJ99] John Weiner, Vanderlei S. Bagnato, Sergio Zilio, and Paul S. Julienne. Experiments and theory in cold and ultracold collisions. Reviews of Modern Physics, 71(1):1–85, January 1999. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.71.1;
http://rmp.aps.org/abstract/RMP/v71/i1/p1_1.

Weissman:1993:WSG

[Wei93] M. B. Weissman. What is a spin glass? A glimpse via mesoscopic noise. Reviews of Modern Physics, 65(3):829–839, July 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.65.829;   http://rmp.aps.org/abstract/
RMP/v65/i3/p829_1.

Weiss:1999:EGR

[Wei99a] Rainer Weiss. Erratum: Gravitational radiation [Rev. Mod. Phys. 71, S187 (1999)]. Reviews of Modern Physics, 71 (4):1247, July 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.1247;   http://
rmp.aps.org/abstract/RMP/v71/i4/p1247_1. See [Wei99b].

Weiss:1999:GR

[Wei99b] Rainer Weiss. Gravitational radiation. Reviews of Modern Physics, 71(2):S187–S196, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.71.S187;   http:
//rmp.aps.org/abstract/RMP/v71/i2/pS187_1. See erratum

[Wei99a].

Walker:1997:SEO

[WH97] Thad G. Walker and William Happer. Spin-exchange optical pumping of noble-gas nuclei. Reviews of Modern Physics, 69(2):629–642, April 1997. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.69.629;   http://
rmp.aps.org/abstract/RMP/v69/i2/p629_1.

Wilczek:1999:QFT

[Wil99] Frank Wilczek. Quantum field theory. Reviews of Modern Physics, 71(2):S85–S95, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.71.S85;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS85_1.

Winckler:1992:CEE

[Win92] John R. Winckler. Controlled experiments in the Earth’s magnetosphere with artificial electron beams. Reviews of Modern Physics, 64(3):859–871, July 1992. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:

//link.aps.org/doi/10.1103/RevModPhys.64.859;   http:/
/rmp.aps.org/abstract/RMP/v64/i3/p859_1.

Wallerstein:1997:SES

[WIP^{+}97] George Wallerstein, Icko Iben, Jr., Peter Parker, Ann Merchant Boesgaard, Gerald M. Hale, Arthur E. Champagne, Charles A. Barnes, Franz Käppeler, Verne V. Smith, Robert D. Hoffman, Frank X. Timmes, Chris Sneden, Richard N. Boyd, Bradley S. Meyer, and David L. Lambert. Synthesis of the elements in stars: forty years of progress. Reviews of Modern Physics, 69(4):995–1084, October 1997. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.69.995;
 http://rmp.aps.org/abstract/RMP/v69/i4/p995_1.

Wiringa:1993:DNS

[Wir93] R. B. Wiringa. From deuterons to neutron stars: variations in nuclear many-body theory. Reviews of Modern Physics, 65 (1):231–242, January 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http:/

/link.aps.org/doi/10.1103/RevModPhys.65.231;   http://
rmp.aps.org/abstract/RMP/v65/i1/p231_1.

Witten:1998:PSG

[Wit98] T. A. Witten. Polymer solutions: a geometric introduction. Reviews of Modern Physics, 70(4):1531–1544, October 1998. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.70.1531;   http://rmp.aps.org/abstract/RMP/
v70/i4/p1531_1.

Witten:1999:ISC

[Wit99] Thomas A. Witten. Insights from soft condensed matter. Reviews of Modern Physics, 71(2):S367–S373, March 1999. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S367;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS367_1.

Wolfenstein:1999:NP

[Wol99] L. Wolfenstein. Neutrino physics. Reviews of Modern Physics, 71(2):S140–S144, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.71.S140;   http://
rmp.aps.org/abstract/RMP/v71/i2/pS140_1.

Wieman:1999:ACT

[WPW99] Carl E. Wieman, David E. Pritchard, and David J. Wineland. Atom cooling, trapping, and quantum manipulation. Reviews of Modern Physics, 71(2):S253–S262, March 1999. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.71.S253;   http://rmp.aps.org/abstract/RMP/
v71/i2/pS253_1.

Watkin:1993:SML

[WRB93] Timothy L. H. Watkin, Albrecht Rau, and Michael Biehl. The statistical mechanics of learning a rule. Reviews of Modern Physics, 65(2):499–556, April 1993. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.65.499;
 http://rmp.aps.org/abstract/RMP/v65/i2/p499_1.

Wu:1992:KTS

[Wu92] F. Y. Wu. Knot theory and statistical mechanics. Reviews of Modern Physics, 64(4):1099–1131, October 1992. CO-DEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (elec-

tronic), 1539-0756. URL http://link.aps.org/doi/10.1103/
RevModPhys.64.1099;   http://rmp.aps.org/abstract/RMP/

v64/i4/p1099_1. See erratum [Wu93].

Wu:1993:EKT

[Wu93] F. Y. Wu. Erratum: Knot theory and statistical mechanics. Reviews of Modern Physics, 65(2):577, April 1993. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-

0756. URL http://link.aps.org/doi/10.1103/RevModPhys.
65.577;   http://rmp.aps.org/abstract/RMP/v65/i2/p577_

1. See [Wu92].

Winstein:1993:SDC

[WW93] Bruce Winstein and Lincoln Wolfenstein. The search for direct CP violation. Reviews of Modern Physics, 65(4):1113– 1147, October 1993. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756. URL http://

link.aps.org/doi/10.1103/RevModPhys.65.1113;   http://
rmp.aps.org/abstract/RMP/v65/i4/p1113_1.

Yarkony:1996:DCI

[Yar96] David R. Yarkony. Diabolical conical intersections. Reviews of Modern Physics, 68(4):985–1013, October 1996. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.68.985;
 http://rmp.aps.org/abstract/RMP/v68/i4/p985_1.

Yu:1995:SAM

[YB95] Zuliang Yu and Siegfried Boseck. Scanning acoustic microscopy and its applications to material characterization. Reviews of Modern Physics, 67(4):863–891, October 1995. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.67.863;
 http://rmp.aps.org/abstract/RMP/v67/i4/p863_1.

Yeshurun:1996:MRH

[YMS96] Y. Yeshurun, A. P. Malozemoff, and A. Shaulov. Magnetic relaxation in high-temperature superconductors. Reviews of Modern Physics, 68(3):911–949, July 1996. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.68.911;
 http://rmp.aps.org/abstract/RMP/v68/i3/p911_1.

Zeilinger:1999:EFQ

[Zei99] Anton Zeilinger. Experiment and the foundations of quantum physics. Reviews of Modern Physics, 71(2):S288–S297, March 1999. CODEN RMPHAT. ISSN 0034-6861 (print), 1538-4527

(electronic), 1539-0756. URL http://link.aps.org/doi/10.
1103/RevModPhys.71.S288;   http://rmp.aps.org/abstract/
RMP/v71/i2/pS288_1.

Zhang:1990:CST

[ZFG90] Wei-Min Zhang, Da Hsuan Feng, and Robert Gilmore. Coherent states: Theory and some applications. Reviews of Modern Physics, 62(4):867–927, October 1990. CODEN RM-PHAT. ISSN 0034-6861 (print), 1538-4527 (electronic), 1539-0756.

URL http://link.aps.org/doi/10.1103/RevModPhys.62.867;
 http://rmp.aps.org/abstract/RMP/v62/i4/p867_1.